### v1

# Discovery and Analysis of Afterglows from Poorly Localised GRBs with the Gravitational-wave Optical Transient Observer (GOTO) All-sky Survey

Abstract
Gamma-ray bursts (GRBs), particularly those detected by wide-field instruments such as the Fermi/GBM, pose a challenge for optical follow-up due to their large initial localisation regions, leaving many GRBs without identified afterglows. The Gravitational-wave Optical Transient Observer (GOTO), with its wide field of view, dual-site coverage, and robotic rapid-response capability, bridges this gap by rapidly identifying and localising afterglows from alerts issued by space-based facilities including Fermi, SVOM, Swift, and the EP, providing early optical positions for coordinated multi-wavelength follow-up. In this paper, we present optical afterglow localisation and multi-band follow-up of seven Fermi/GBM and MAXI/GSC triggered long GRBs (240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B) discovered by GOTO in 2024. Spectroscopy for six GRBs (no spectroscopic data for GRB 241002B) with VLT/X-shooter and GTC/OSIRIS yields precise redshifts spanning $z\approx 0.40$–3.16 and absorption-line diagnostics of host and intervening systems. Radio detections for four events confirm the presence of long-lived synchrotron emission. Prompt-emission analysis with Fermi and MAXI data reveals a spectrally hard population, with two bursts lying $3\sigma$ above the Amati relation. Although their optical afterglows resemble those of typical long GRBs, the prompt spectra are consistently harder than the long-GRB average. Consistent modelling of six GOTO-discovered GRB afterglows yields jet half-opening angles of a few degrees and beaming-corrected kinetic energies $E_{\rm jet}\sim 10^{51}$$10^{52}$ erg, consistent with the canonical long-GRB population. These findings suggest that optical discovery of poorly localised GRBs may be subject to observational biases favouring luminous events with high spectral peak energy ($E_{\rm p}$), while also providing insight into jet microphysics and central engine diversity.
keywords:
transients: gamma-ray bursts – gamma-ray bursts: general – gamma-ray burst: individual: GRB 240122A, GRB 240225B, GRB 240619A, GRB 240910A, GRB 240916A, GRB 241002B, GRB 241228B – techniques: photometric – techniques: spectroscopic.
pubyear: 2025pagerange: Discovery and Analysis of Afterglows from Poorly Localised GRBs with the Gravitational-wave Optical Transient Observer (GOTO) All-sky SurveyAffiliations
1 Introduction
The study of gamma-ray bursts (GRBs) has advanced significantly since their discovery in the 1960s (Klebesadel et al., 1973; Strong et al., 1974), driven by a combination of dedicated space-based surveys and ground-based follow-up observations. From confirming their cosmological origins (Meegan et al., 1992; Costa et al., 1997; Metzger et al., 1997; van Paradijs et al., 1997) to uncovering possible progenitors for long and short GRBs (Mazets et al., 1981; Kouveliotou et al., 1993; Woosley & Bloom, 2006; Zhang et al., 2012; Abbott et al., 2017b), these high-energy events have now been recognised as one of the most powerful explosions in the universe (Mészáros, 2013; Kumar & Zhang, 2015; Levan et al., 2016; LHAASO Collaboration et al., 2023). Observations across the electromagnetic spectrum have not only enhanced our understanding of GRB mechanisms but also established their ability to probe the distant universe (Fiore, 2001; Tanvir et al., 2009; Petitjean & Vergani, 2011; Saccardi et al., 2023, 2025) and constrain cosmological parameters (Amati & Della Valle, 2013; Demianski et al., 2017; Luongo & Muccino, 2021; Moresco et al., 2022), marking them as invaluable tools in modern astrophysics.
[…]
The origin of GRBs and the understanding of the underlying physics can be probed using multi-wavelength afterglow observations (see Miceli & Nava 2022). Synchrotron emission from relativistic jets that interact with the circumburst medium encodes information about the jet geometry, ambient density, and microphysical parameters (Sari et al., 1998; Granot & Sari, 2002; Panaitescu & Kumar, 2002), see also Zhang et al. (2024). Afterglow light curve features such as jet breaks, cooling breaks, and chromatic evolution offer insights into jet collimation and energy structure (Rhoads, 1999; Sari & Piran, 1999). Theoretical models provide further context: relativistic jet propagation in collapsars was first explored via simulations (Aloy et al., 2000), and more recent 3D magnetorotational core-collapse models further demonstrate jet collimation and dynamics in magnetised environments (Obergaulinger & Aloy, 2021). Complementary hydrodynamical studies examine jet–cocoon mixing and structured outflow morphologies (Gottlieb et al., 2020).
Early optical and multi-wavelength follow-up has revealed a broad diversity in afterglow behaviours, including evidence for reverse shocks (Zhang & Kobayashi, 2005; Mundell et al., 2007; Mimica et al., 2009, 2010; Laskar et al., 2013; Yi et al., 2020), energy injection episodes (Björnsson et al., 2004; Zhang et al., 2006; Laskar et al., 2015), and structured jets (Lamb & Kobayashi, 2017; Beniamini et al., 2020; Oganesyan et al., 2020). Combined with X-ray and radio data, optical observations enable comprehensive modelling of afterglows, shedding light on the energetics and structure of GRB jets (Margutti et al., 2013). However, a persistent challenge in GRB afterglow detection and follow-up arises from the poor initial localisation provided by wide-field gamma-ray monitors such as Fermi Gamma-ray Burst Monitor (GBM; Meegan et al. 2009). With the highest GRB detection rate and strong sensitivity to prompt gamma-ray emission, Fermi/GBM enables detailed temporal and spectral studies (Meegan et al., 2009; von Kienlin et al., 2020), but typically provides localisation uncertainties spanning several square degrees. In contrast, missions like Neil Gehrels Swift Observatory (Swift hereafter; Gehrels et al. 2004), Einstein Probe (EP; Yuan et al. 2015, 2022), and Space-based multi-band astronomical Variable Objects Monitor (SVOM; Wei et al. 2016) offer arcsecond–arcminute level localisations but detect comparatively fewer bursts. Recovering counterparts to these poorly localised GRBs provides a valuable opportunity to expand our understanding of GRB diversity and reduce selection biases. However, the large error regions often exceed the field of view (FoV) of conventional optical telescopes, complicating timely afterglow identification. Additionally, the intrinsic faintness of some afterglows (Liang et al., 2007; Dereli et al., 2017) and circumburst extinction (Savaglio & Fall, 2004; Schulze et al., 2011) can further hinder follow-up. Without alternative localisation strategies, a significant fraction of GRBs, particularly those detected by Fermi/GBM, remain uncharacterised, limiting our ability to probe jet physics, energetics, and progenitor properties.
Wide-field optical instruments, such as the Gravitational-wave Optical Transient Observer (GOTO [^2] https://goto-observatory.org/; Dyer et al. 2020; Steeghs et al. 2022), have emerged as powerful tools to address this gap. GOTO, comprising 32 robotic telescopes across two sites, Roque de los Muchachos Observatory (La Palma, Canary Islands) and Siding Spring Observatory (New South Wales, Australia), enables near-continuous coverage of both hemispheres (Dyer et al., 2024). Unlike traditional follow-up facilities, the GOTO instruments can “tile” the large error regions of GBM-like triggers in near real-time, providing a complementary discovery channel to narrow-field missions and helping to overcome localisation-driven selection effects. Its fast-response capabilities and wide FoV make it well-suited to bridging the gap between gamma-ray detection and multi-wavelength characterisation, particularly for poorly localised but scientifically valuable GRBs.
In this work, we present a systematic study of seven poorly localised LGRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B, whose optical afterglows were discovered by GOTO in response to alerts from Fermi/GBM and the Monitor of All-sky X-ray Image (MAXI) Gas Slit Camera (GSC; Matsuoka et al. 2009; Mihara et al. 2011), with localisation uncertainties ranging from a few arcminutes to several square degrees in radius. The structure of this paper is as follows. Section 2 provides an overview of GOTO, outlining its observational strategy and summarising its past observation records. Section 3 introduces the sample and describes the discovery of their afterglows in terms of localisation coverage and optical afterglow detections. Section 4 presents the multi-wavelength (X-ray to radio) follow-up observations of the afterglows. Section 5 details the prompt gamma-ray analyses and properties of the GRBs in our sample, along with comparisons to other GRBs. Section 6 discusses the multi-wavelength afterglow properties and compares them with literature data for other GRBs, including optical spectroscopic analyses of the afterglows and precise redshift estimates. Section 7 presents the afterglow modelling of six GRBs from our sample using afterglowpy and Bayesian inference with dynesty nested sampling. Finally, Section 8 summarises our findings and presents the main conclusions of this study.
All magnitudes reported in this work are given in the AB photometric system. We define $T_{0}$ as the trigger time reported by the detecting satellite, which serves as the reference epoch for our temporal analysis. For GRBs 240122A and 240225B this corresponds to the MAXI/GSC trigger, while for the remaining five events it corresponds to the Fermi/GBM trigger.
2 GOTO and Its Approach to GRB Counterpart Searches
2.1 GOTO Overview
The Gravitational-wave Optical Transient Observer (GOTO; Dyer et al., 2020; Steeghs et al., 2022; Dyer et al., 2024) is a global network of 32 robotic unit telescopes (UTs) distributed over two sites, with two domes at each site: the Roque de los Muchachos Observatory on La Palma, Canary Islands, and the Siding Spring Observatory in New South Wales, Australia (see Figure 1). This configuration enables near-continuous monitoring of both the northern and southern skies.
Each site hosts two independent mounts, each carrying eight UTs with a combined instantaneous FoV of $\approx 44\,deg^{2}$. Across the 16 UTs at both sites, GOTO achieves a total FoV of $\approx 88\,deg^{2}$, making it well-suited to search for optical counterparts of poorly localised transients such as gravitational-wave events (e.g., Gompertz et al., 2020), GRBs detected by facilities such as Fermi, SVOM and EP (e.g., Mong et al., 2021; Belkin et al., 2024), as well as other fast and exotic transients including rapidly-evolving supernovae and tidal disruption events. The telescopes are equipped with ON Semiconductor KAF-50100 CCDs, which provide broad sensitivity across the optical range, with the deployed Baader filters setting the effective bandpass. In survey mode, a wide $L$-band (400700400-700 nm) encompassing the Sloan gri filters is used, providing sensitivity to a wide range of transients and maximising discovery potential.
In “responsive” mode, the GOTO instruments autonomously observe large sky regions associated with poorly localised transient events, such as GW events, GRBs, and high-energy neutrino alerts, to search for their optical counterparts. The exposure time and cadence in this mode are adapted to the nature of the event and vary accordingly across different source types. This study focuses specifically on the discovery of optical afterglows from poorly localised GRBs using GOTO. The following section outlines the observational strategies employed by GOTO to identify and confirm these afterglows.
2.2 Follow-up Strategies to Discover GRBs’ Optical Afterglows
Figure 1: The full configuration of the GOTO telescope network in April 2023, comprising 32 robotic unit telescopes distributed across four domes, two domes at each of the two sites. Top: GOTO-N, located at the Observatorio del Roque de los Muchachos on La Palma, comprising GOTO-1 (left) and GOTO-2 (right). Bottom: GOTO-S, hosted at Siding Spring Observatory in Australia, consisting of GOTO-3 (left) and GOTO-4 (right). Figure credit: Dyer et al. 2024.
In responsive mode, if triggered by a GRB alert, GOTO pauses its survey operations to target the localisation region. Follow-up strategies are tailored based on the source of the trigger and the localisation uncertainty, as described below and illustrated in the flowchart in Figure 2.
Swift/BAT, SVOM/ECLAIRs, and EP triggers: Swift, SVOM/ECLAIRs, and EP GRB detections generally come with precise localisation (arcsecs to arcmins), far smaller than a single GOTO tile, and which can be easily covered by other observatories with a lower FoV. Therefore, the primary reason for GOTO to follow up these events is to take advantage of its fast, robotic nature to get rapid coverage of the localisation region immediately following the trigger, in order to capture any optical afterglow while it is still bright and young. For these triggers, up to five tiles on the GOTO survey grid are permitted to be selected; however, given the well-localised nature of these sources, the search region is almost always within a single tile (more than one tile is allowed to be selected for rare cases where a source falls within the overlapping region on the edge of multiple tiles). Two observations are scheduled for each tile, spaced one hour apart, each using the standard set of $4\times 90$ s exposures, which typically reaches a depth of $\sim 19.8$ AB mag in GOTO $L$-band. However, these targets are only valid in the GOTO scheduling queue for the first two hours after the trigger time. This ensures that rapid observations will be taken if any of the GOTO telescopes are available immediately after the trigger; however, after the two-hour window, GOTO’s usefulness is lessened, and any observations are left to other observatories.
Figure 2: A summary of the GOTO GRB follow-up strategy.
[…]
2.3 Identifying the GRB Optical Counterpart
For each GOTO observing sequence, whether taken in responsive or survey mode, images are processed in near real-time by the GOTO pipeline (Lyman et al., in prep.), which includes calibration, astrometric solution, and difference imaging against archival deep GOTO templates. Transient candidates are then ranked by a machine-learning classifier (Killestein et al., 2021), cross-matched to contextual catalogues (e.g., SDSS York et al. 2000; Pan-STARRS Chambers et al. 2016), and Solar System ephemerides, and subsequently passed through automated real/bogus and contextual filters.
The Burst Advocate (BA) monitors GRB alerts, confirms that GOTO follow-up has been executed, and initiates candidate vetting in the GOTO marshall. Promising sources are inspected by the working group, including checks against archival imaging and forced photometry, before being promoted for group review. A candidate is classified as the optical afterglow counterpart if it satisfies the following:
  1. 1.
    Spatial consistency: positionally coincident with the high-probability GRB localisation region, with a point-like PSF and no association with known artefacts or moving objects (minor-planet checks performed). If present, the location relative to a plausible host galaxy is also considered.
  2. 2.
    Temporal behaviour: evidence of fading between successive GOTO epochs, or a later non-detection deeper than the initial detection; where possible, the decline should be consistent with power-law afterglow behaviour.
  3. 3.
    Contextual screening: absence of a persistent source in archival templates; no counterpart in variable-star catalogues, and not coincident with a known AGN nucleus.
  4. 4.
    Multi-wavelength corroboration: spatial consistency with a Swift/XRT source strengthens the association, but is not required.
Candidates satisfying criteria (i)–(iii) are promoted to the transient stream and considered GRB afterglow counterparts, while those also fulfilling (iv) are prioritised for rapid spectroscopy and additional ToO follow-up. Confirmed counterparts are reported in General Coordinates Network (GCN [^3] https://gcn.nasa.gov/circulars) circulars and logged in the Transient Name Server (TNS [^4] https://www.wis-tns.org/).
2.4 Observational Performance and Discovery Statistics
GOTO’s GRB follow-up programme has developed from its prototype stage into a fully operational dual-site facility (Dyer et al., 2024), delivering a series of notable discoveries. The prototype system (GOTO-4), comprising four UTs at the La Palma site, achieved first light in June 2017 and was officially inaugurated in July 2017, initiating routine operations (Steeghs et al., 2022). During this early phase, GOTO secured its first GRB optical afterglow detection with GRB 171205A, associated with SN 2017iuk (Steeghs et al., 2017; Izzo et al., 2019). In the initial 3 years between June 2017 and June 2020, the GOTO-4 system responded to 77 Fermi/GBM and 29 Swift/BAT triggers (see Mong et al., 2021; Steeghs et al., 2022), demonstrating the scientific potential of the facility even in its prototype configuration.
With full deployment at both sites, as of 31 December 2024 (with GRBs thereafter observed to be included in a future study), GOTO has conducted follow-up observations for over 257 Fermi -, 43 Swift -, 28 EP-, and 7 GECAM-triggered events. No targeted observations were conducted for SVOM events during this period, as GOTO began following up SVOM triggers in 2025, coinciding with the scheduled start of SVOM’s science operations in February 2025. On average, GOTO’s first targeted observation latency was approximately 11.2 hr, with response times ranging from 270 s–69.77 hr post-trigger. The average latency for the last observation in each follow-up series was 26.36 hr, while the alert latency averaged 4.8 hr, underscoring the challenges in achieving timely and efficient follow-ups.
Table 1: Summary of the LGRBs analysed in this work. For each burst, we list the high-energy and optical afterglow discoverers, the time from the high-energy trigger to the optical discovery ($T-T_{0}$), the GOTO internal afterglow name, J2000 coordinates, discovery $L$-band magnitude, Galactic extinction E(BV)E(B-V), and spectroscopic redshift.
GRB High-energy Optical Discovery GOTO RA Dec Discovery $E(B-V)^{b}$ Redshiftc
discoverer discoverer $T-T_{0}$ (hrs)a internal name (h:m:s) (::′′) $L$-band mag (mag)
240122A MAXI/GSC GOTO-S 0.73 GOTO24eu 06:12:12.91 -19:08:38.81 $17.58\pm 0.04$ 0.0651 $3.1634\pm 0.0003$
240225B MAXI/GSC GOTO-N 1.50 GOTO24tz 08:33:26.67 +27:04:32.71 $17.12\pm 0.04$ 0.0354 $0.9462\pm 0.0002$
240619A Fermi/GBM GOTO-S 4.69 GOTO24cvn 10:49:34.70 +17:16:58.07 $17.17\pm 0.17$ 0.0253 $0.3960\pm 0.0001$
240910A Fermi/GBM GOTO-S 9.43 GOTO24fvl 01:36:23.45 -00:12:17.86 $19.33\pm 0.13$ 0.0247 $1.4605\pm 0.0007$
240916A Fermi/GBM GOTO-S 7.73 GOTO24fzn 15:43:39.23 -07:45:53.21 $17.80\pm 0.06$ 0.1359 $2.6100\pm 0.0002$
241002B Fermi/GBM GOTO-S 3.05 GOTO24gpc 21:53:16.56 -58:56:51.98 $19.53\pm 0.09$ 0.0268
241228B Fermi/GBM GOTO-N 0.32 GOTO24jmz 08:31:05.46 +06:50:54.07 $14.54\pm 0.01$ 0.0290 $2.6745\pm 0.0004$
  • a
    For MAXI GRBs, $T_{0}$ denotes the MAXI/GSC trigger time; for Fermi GRBs, $T_{0}$ denotes the Fermi/GBM trigger time.
  • b
    Galactic extinction values are estimated following recalibrated dust maps of Schlafly & Finkbeiner (2011).
  • c
    see Section 6.3.
The GOTO collaboration has reported nearly 80 GCN circulars to date based on GOTO observations, including the detection and upper-limit constraints of GRB afterglows. To date, nearly 28 successful detections have been reported, including GRB 230818A, detected 4.43 min after its trigger (Gompertz et al., 2023c). In addition, GOTO has provided numerous upper-limit constraints and contributed to serendipitous discoveries like orphan afterglow AT2023lcr (Gompertz et al., 2023b), with early-phase observations aiding in the refinement of transient properties (Martin-Carrillo et al., in prep.). GOTO’s follow-up capabilities continue to evolve in response to operational experience and scientific objectives.
GRB 230911A was the first LGRB for which GOTO discovered an optical afterglow (Belkin et al., 2023); details are published in Belkin et al. (2024). After this first case, in 2024, GOTO discovered optical afterglows of 2 MAXI-triggered (GRB 240122A and GRB 240225B) serendipitously and 5 Fermi-triggered (GRBs 240619A, 240910A, 240916A, 241002B and 241228B) LGRBs in responsive mode, which are studied in detail in the present work. The details of these 7 GRBs in our sample are discussed in the following section. In addition to these LGRBs, within 2024 itself, GOTO also identified the optical afterglow of the SGRB 241105A (Julakanti et al., 2024b), which was localised by Fermi/GBM with an uncertainty of $\sim 4\,\mathrm{deg}$ (Fermi GBM Team, 2024e). GOTO rapidly responded to the trigger, tiling 277.9 square degrees within the 90% GBM localisation region and covering $\sim 84\%$ of the total probability within 1.6 hr. The afterglow was discovered at $L\sim 17.2$ mag, later confirmed through multi-wavelength follow-up and spectroscopy to lie at a redshift of $z=2.681$ (Izzo et al., 2024). Although this event is not part of the core LGRB sample analysed here, it highlights GOTO’s capabilities to detect optical afterglows from both long and short GRBs, even under challenging localisation conditions. A detailed analysis of GRB 241105A is presented in a separate paper by Dimple et al. (2025).
3 The GRB sample
Our sample comprises seven LGRBs whose optical afterglows were discovered by GOTO in 2024: two MAXI–triggered events (GRBs 240122A and 240225B) identified serendipitously, and five Fermi–triggered events (GRBs 240619A, 240910A, 240916A, 241002B, and 241228B) detected in responsive mode. This section begins by discussing the prompt high-energy triggers and observations of these seven GRBs. The basic properties of the GRBs in our sample are listed in Table 1. The detection circumstances provide essential context — in particular, the trigger times, localisation accuracy, and alert distribution — that directly influenced GOTO’s follow-up strategy, as discussed above in Section 2.2. The subsequent subsections describe the localisation, follow-up coverage, and an observational summary of the seven LGRBs in our sample.
3.1 High-Energy Triggers
The GOTO follow-up campaigns for GRBs in our sample were initiated by triggers from the MAXI/GSC (Matsuoka et al., 2009; Mihara et al., 2011) and the Fermi/GBM (Meegan et al., 2009). We briefly summarise below the specific GRBs that prompted these observations. Figure 3 shows the 90% containment localisation regions provided by the triggering satellites.
3.1.1 MAXI/GSC GRBs 240122A and 240225B
GRB 240122A was detected by the MAXI/GSC on 2024-01-22 at 10:28:03 UT in the 2–10 keV range (Negoro et al., 2024). The burst was localized to RA = 06h11m18s and Dec = -190151′′ (J2000), with an uncertainty of 30 arcmin. This event was detected solely by MAXI/GSC, with no additional high-energy instruments reporting a detection. Due to the large localisation uncertainty, it was not followed up by narrow-field optical instruments but represented a good candidate for wide-field facilities such as GOTO.
MAXI/GSC triggered GRB 240225B on 2024-02-25 at 20:15:46 UT in the 4–10 keV range (Nakajima et al., 2024). The burst was localized to RA = 08h33m49s and Dec = +272913′′ (J2000), with a statistical 90% confidence level elliptical error region, where the semi-major and semi-minor axes have radii of 0.13° and 0.11°, respectively (see Figure 3). The burst was also observed by several other high-energy instruments, including AstroSat’s Cadmium Zinc Telluride Imager (CZTI) in the 20–200 keV band, and the CsI anticoincidence detectors in the 100–500 keV band (Joshi et al., 2024). Additional detections include INTEGRAL SPI–Anti-Coincidence Shield (SPI–ACS) in energies $\gtrsim$80 keV, the CALET Gamma-ray Burst Monitor (CGBM) in the 100–500 keV range and coverage up to 40–1000 keV (Kawakubo et al., 2024), Konus-Wind up to $\sim$3 MeV (Frederiks et al., 2024), and the Glowbug gamma-ray telescope in the 10–10000 keV range (Cheung et al., 2024).
3.1.2 Fermi/GBM GRBs 240619A, 240910A, 240916A, 241002B, and 241228B
GRB 240619A triggered the Fermi/GBM on 2024-06-19 at 03:43:31 UT in the 50–300 keV band (Fermi GBM Team, 2024a; Dalessi et al., 2024). The burst was localised to RA = 10h48m00s and Dec = +171800′′ (J2000), with a statistical uncertainty of 1.6°, which is circularised 90% containment radius $R_{\mathrm{err,90}}$ in degrees (see Figure 3 and column two of Table 2). In addition to Fermi/GBM, the burst was also detected by the CALET CGBM in the 40–1000 keV band (Torii et al., 2024), Konus–Wind up to $\sim$10 MeV (Svinkin et al., 2024), and the 1U-sized CubeSat GRBAlpha in the 70–890 keV band (Dafcikova et al., 2024a).
GRB 240910A triggered the Fermi/GBM on 2024-09-10 at 04:00:44 UT (Fermi GBM Team, 2024b). The event was localised to RA = 01h00m00s and Dec = +043000′′ (J2000), with a statistical uncertainty of 4.5°. This burst was also picked up by the SVOM/GRM operating in the 15 keV–5 MeV range (SVOM/GRM Team et al., 2024), as well as by the 1U CubeSat GRBAlpha (Ripa et al., 2024).
GRB 240916A, detected by the Fermi/GBM at 01:22:56 UT on 2024-09-16 (Fermi GBM Team, 2024c; Roberts et al., 2024a), was localised to RA = 15h32m00s and Dec = -070500′′ (J2000) with a statistical uncertainty of 1.2°. Additional high-energy observations were made by the INTEGRAL/SPI–ACS (Pawar, 2024) and by the 1U CubeSat GRBAlpha (Dafcikova et al., 2024b).
GRB 241002B triggered the Fermi/GBM at 06:14:18.76 UT on 2024-10-02 (Fermi GBM Team, 2024d; Roberts et al., 2024b). The burst was localised to RA = 22h15m00s and Dec = -641700′′ (J2000), with a statistical uncertainty of 3.7°. It was also observed by the Swift Burst Alert Telescope – Gamma-ray Urgent Archiver for Novel Opportunities (Swift/BAT–GUANO; DeLaunay et al. 2024a).
GRB 241228B triggered the Fermi/GBM at 04:13:05.39 UT on 2024-12-28 (Fermi GBM Team, 2024f; Scotton et al., 2024). The burst was localised to RA = 08h08m00s and Dec = +140000′′ (J2000), with a statistical uncertainty of 1.7°. It was also detected by the Swift/BAT–NITRATES system (DeLaunay et al., 2024b), with the position consistent with the GBM localisation. In addition, the Fermi Large Area Telescope (LAT) observed high-energy emission (>100>100 MeV) from this burst, including a 16 GeV photon detected 31 s after the trigger (Di Lalla et al., 2025).
These prompt high-energy detections by Fermi/GBM provided the initial localisation constraints and trigger alerts that enabled rapid optical follow-up by GOTO. The diversity in localisation uncertainties influenced the choice of tiling patterns and observing cadences. In the following section, we present the optical afterglow localisations and follow-up coverage, highlighting how the prompt trigger information shaped the subsequent GOTO observations.
3.2 Localisation Coverage and Optical Afterglow Discoveries
Figure 3 provides an overview of the sky localisation and follow-up coverage for the GRBs analysed in this study (GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B), based on GOTO observations obtained within the first 24 h post-trigger. The dark and light grey contours represent the $1\sigma$ and $2\sigma$ localisation regions extracted from the HEALPix (Górski et al., 2005) probability skymaps. Light blue shading denotes the GOTO fields observed within 10 hr of each burst, demonstrating the system’s wide-area and rapid-response capabilities. Cyan stars mark the locations of confirmed optical afterglows, showing that GOTO’s coverage either enclosed or closely bordered the true source positions in all cases. In addition, the finding charts highlighting the confirmed afterglows, derived from GOTO observations, are presented in panels (a)–(g) of Figure 4. For context, each GOTO cutout is paired with a corresponding image from the DESI Legacy Survey (LS) DR10 covering the same $\sim 6.3^{\prime}\times 6.3^{\prime}$ FoV; for GRB 240122A, a Pan-STARRS DR1 image is shown instead.
3.2.1 MAXI/GSC GRBs 240122A and 240225B: arcminute-scale localisation
GOTO does not follow up MAXI triggers in responsive mode, as MAXI alerts are not distributed in a machine-readable format that can be ingested automatically by the sentinel. Instead, these fields are only covered serendipitously in survey mode. In Figure 3, the top row shows the two events detected by MAXI/GSC, GRB 240122A and GRB 240225B, each displayed within a $3^{\circ}\times 3^{\circ}$ FoV. The red shaded regions indicate the approximate localisation areas from MAXI/GSC reports. For GRB 240122A, the afterglow is well centred within the localisation and fully encompassed by the GOTO field. In contrast, for GRB 240225B, the GOTO tiling intersected the elongated error region, providing timely coverage that included the eventual afterglow position.
GRB 240122A
GOTO-S serendipitously observed the localisation region of GRB 240122A during its routine all-sky survey on 2024-01-22, discovering the optical afterglow (GOTO24eu/AT2024apy) at J2000 coordinates RA = 06h12m12s.91 and Dec = -190838′′.81. The afterglow was detected at 11:11:43 UT (T0+43.68 min) with an $L$-band magnitude of $17.06\pm 0.04$ mag (Kumar et al., 2024b), see also Table 1 for details. As shown in Figure 3 (top-left), the GOTO FoV comfortably covers the compact MAXI/GSC localisation, placing the afterglow well within the imaged area. This case demonstrates GOTO’s ability to capture transient counterparts during its high-cadence survey mode, even without a targeted trigger.
Table 2: Summary of GOTO’s coverage of the Fermi/GBM GRBs in our sample. For each event, we list the Fermi/GBM-reported 90% localisation uncertainty ($R_{\mathrm{err,90}}$), number of images, sky coverage area, enclosed probability, and mean $5\sigma$ $L$-band limiting magnitude.
    GRB     $R_{\mathrm{err,90}}$     No. of     Coverage     Prob.     Mean $5\sigma$
         (deg)     images     (deg2)     enclosed (%)     depth (mag)
    240619A     1.6     56     151.5     85.7     18.6
    240910A     4.5     191     295.3     90.3     20.0
    240916A     1.2     41     178.6     78.3     19.1
    241002B     3.7     58     273.1     84.9     20.3
    241228B     1.7     165     214.0     89.9     19.8
GRB 240225B
Similarly, GOTO-N serendipitously covered the field of GRB 240225B and discovered its optical afterglow (GOTO24tz/AT2024dgu) at J2000 coordinates RA = 08h33m26s.67 and Dec = +270432′′.71. The counterpart was first detected on 2024-02-25 at 21:45:51 UT (T0+1.50 h) at $17.12\pm 0.04$ mag ($L$-band) and was last detected the following night at 22:10:38 UT (T0+25.91 h) at $19.69\pm 0.18$ mag (Gompertz et al., 2024a). As shown in Figure 3 (top-right), the GOTO tiling intersected the elongated MAXI/GSC localisation, with the afterglow located near the centre of the observed field.
Figure 3: GOTO coverage of each of the GRBs in the sample. The first two plots denote the 90% containment MAXI/GSC localisations (red) in a $3\degree\times 3\degree$ field. The localisation areas are generated based on information from their discovery GCNs. The following five plots show Fermi/GBM localisations (grey) and the 1 and 2 $\sigma$ contours from their respective HEALPix skymaps in a $20\degree\times 20\degree$ field. In all plots, the 2D footprint of GOTO images taken in the first 10 hr post-trigger that overlap the localisations are shown in light blue. The corresponding afterglow positions are marked with a cyan star.
GRB 240122A    (a) GRB 240225B    (b)
Refer to caption Refer to caption Refer to caption Refer to caption
GRB 240619A    (c) GRB 240910A    (d)
Refer to caption Refer to caption Refer to caption Refer to caption
GRB 240916A    (e) GRB 241002B    (f)
Refer to caption Refer to caption Refer to caption Refer to caption
GRB 241228B    (g)
Refer to caption Refer to caption
Figure 4: Finding charts of GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B in the GOTO $L$-band (400–700 nm) observed by GOTO. Each cutout is a $300\times 300$ pixel region centred on the transient, corresponding to a FoV of $\sim 6.3\arcmin\times 6.3\arcmin$ at the GOTO pixel scale of $1.26\arcsec$/pix. For comparison, survey images from the DESI Legacy Survey DR10 are shown (except for GRB 240122A, where a Pan-STARRS DR1 image is used), matched to the same FoV. Details of each object are listed in Table 1.
3.2.2 Fermi/GBM GRBs 240619A, 240910A, 240916A, 241002B, and 241228B: degree-scale localisation
Fermi/GBM triggers are distributed in real time via machine-readable GCN notices, which the GOTO sentinel ingests automatically. This enables fully responsive follow-up, with observations scheduled immediately after the alert is received. Follow-up observations are carried out as soon as observing conditions and visibility constraints permit. The five different panels in Figure 3 correspond to events detected by Fermi/GBM GRBs 240619A, 240910A, 240916A, 241002B, and 241228B, each displayed over a wider $20^{\circ}\times 20^{\circ}$ field. A quantitative summary of GOTO follow-up coverage for the Fermi/GBM GRBs in our sample is provided in Table 2. For each event, we list the GBM localisation uncertainty (expressed as the circularised 90% containment radius in degrees), number of images obtained to cover the 90% GBM localisation region, the total sky area imaged, the fraction of localisation probability enclosed within the observed fields and the mean $5\sigma$ limiting magnitude. The two MAXI/GSC GRBs are not included here, as their compact localisations were fully covered serendipitously by single $\sim 9$ deg2 survey pointings.
GRB 240619A
GOTO-S initiated targeted follow-up observations of GRB 240619A on 2024-06-19 at 08:24:01 UT (T0+4.68 h), continuing until 21:48:35 UT (T0+18.08 h). As illustrated in Figure 3 (second row; left panel), the GOTO tiling successfully overlapped the $1\sigma$ and $2\sigma$ HEALPix contours from the Fermi/GBM localisation, with the afterglow position (blue star) falling within the observed fields obtained in the first 10 hr. The afterglow (GOTO24cvn/AT2024lwv) was identified in these data at J2000 coordinates RA = 10h49m34s.70 and Dec = +171658′′.07, with detections by GOTO-S at 08:24:50 UT (T0+4.69 h) and by GOTO-N at 21:40:50 UT (T0+18.00 h), exhibiting $L$-band magnitudes of $17.17\pm 0.17$ and $18.38\pm 0.09$ mag, respectively (Gompertz et al., 2024b), details are tabulated in Table 1. Although the afterglow was first discovered by GOTO, the position was also serendipitously covered by the ATLAS all-sky survey (Tonry et al., 2018), which provides a forced photometric detection at an earlier epoch. The source is detected in the ATLAS forced photometry data with an $o$-band (560–820 nm) magnitude of $16.24\pm 0.01$ at T0+2.53 hr and $18.72\pm 0.12$ at T0+26.71 hr, retrieved from the ATLAS Forced Photometry Server (Shingles et al., 2021). While ATLAS did not identify the transient in real time, its archival data proved valuable in constraining the early-time brightness and confirming the fading behaviour consistent with an optical afterglow (Gompertz et al., 2024b).
GRB 240910A
GOTO-S began targeted follow-up observations of GRB 240910A on 2024-09-10 at 12:25:28 UT (T0+8.41 h), continuing through to 16:33:57 UT (T0+12.55 h). As shown in Figure 3 (second row; right panel), the GOTO coverage successfully intersected the high-probability localisation contours, with the afterglow position clearly lying within the imaged area. The afterglow (GOTO24fvl/AT2024vfp) was discovered at J2000 coordinates RA = 01h36m23s.45 and Dec = -001217′′.86, with detections spanning from 13:26:31.776 UT (T0+9.43 h) to 16:01:36 UT (T0+12.01 h), yielding $L$-band magnitudes of $19.32\pm 0.13$ and $19.74\pm 0.12$ mag, respectively (Julakanti et al., 2024a).
GRB 240916A
GOTO-S began targeted follow-up observations of GRB 240916A on 2024-09-16 at 09:06:47 UT (T0+7.73 h), continuing until 09:23:05 UT (T0+8.00 h). As seen in Figure 3 (third row; left panel), the GOTO tiling intersected the high-probability regions of the localisation, successfully encompassing the afterglow site. The afterglow (GOTO24fzn/AT2024vlp) was identified at J2000 coordinates RA = 15h43m39s.229 and Dec = -074553′′.22, with a detection at 09:06:47.81 UT (T0+7.73 h) at an $L$-band magnitude of $17.80\pm 0.05$ mag (Gompertz et al., 2024c).
GRB 241002B
GOTO-S conducted targeted follow-up observations of GRB 241002B starting on 2024-10-02 at 09:17:03 UT (T0+3.05 h), concluding at 09:40:00 UT (T0+3.43 h). As shown in Figure 3 (third row; right panel), the observed GOTO fields overlapped the high-probability regions of the Fermi/GBM localisation, with the afterglow position included within the footprint. The afterglow (GOTO24gpc/AT2024xbg) was discovered at J2000 coordinates RA = 21h53m16s.56 and Dec = -585651′′.98, with a detection at 09:17:20 UT (T0+3.05 h) at an $L$-band magnitude of $19.53\pm 0.09$ mag (Kumar et al., 2024c).
GRB 241228B
GOTO-N initiated follow-up observations of GRB 241228B on 2024-12-28 at 04:26:19 UT (T0+0.22 h), continuing through to 23:24:22 UT (T0+19.19 h). As depicted in Figure 3 (bottom-left panel), the GOTO tiling efficiently covered the high-probability localisation region, including the afterglow position. The afterglow (GOTO24jmz/AT2024afgu) was identified by GOTO-N at J2000 coordinates RA = 08h31m05s.46 and Dec = +065054′′.07, with an initial detection at 04:32:24 UT (T0+19.3 min) at an $L$-band magnitude of $14.54\pm 0.01$ mag. Multiple detections followed throughout the observing sequence, with the final GOTO-N detection recorded at 13:00:42 UT (T0+8.79 h) at $19.70\pm 0.10$ mag (Kumar et al., 2024d). Afterglow candidate for GRB 241228B lies on the 94.5% probability contour, formally outside the GBM 90% localisation region. While most GRB afterglows are found within the 90% contour, a small fraction are expected to lie just beyond it, making GRB 241228B a noteworthy case.
Figure 5: First detection times across various observatories for GRBs in our sample. Shown are prompt (red), X-ray afterglow (blue), UV/optical/NIR afterglow (green), and radio afterglow (purple) observations from both space- and ground-based facilities. In all cases, GOTO discovered the optical afterglow following the prompt emission. For GRB 240619A, although ATLAS has the earliest epoch, the afterglow was first discovered by GOTO, and the ATLAS data were serendipitously pre-covered and used in the GOTO discovery report (Gompertz et al., 2024b).
4 Afterglow Follow-up Observations
In addition to the discovery imaging provided by GOTO, we carried out a coordinated programme of multi-wavelength follow-up to characterise the afterglows of our GRB sample. These observations span the X-ray, UV, optical, and radio regimes, enabling us to track the temporal evolution and spectral energy distributions of the counterparts. The combined dataset allows us to constrain the physical properties of the bursts, verify their association with the optical transients identified by GOTO, and provide essential input for modelling their afterglow emission.
Figure 5(a–g) illustrates the timeline of the multi-wavelength follow-up campaigns for all GRBs in the sample, marking the epochs at which each facility recorded its first observation relative to the trigger time. The various phases of the events are colour-coded as follows: X-ray afterglow (blue), UV/optical/NIR afterglow (green), radio afterglow (purple), and prompt emission (red) for completeness. This timeline highlights the wide temporal coverage and the rapid, coordinated response from both ground- and space-based observatories across the electromagnetic spectrum for the GRBs in our sample. More observation details are provided below. A full log of photometric measurements, combining GCN Circular reports and data from this work, is presented in Tables LABEL:tab:phot_data and A2, while Table 3 summarises the spectroscopic campaigns.
4.1 X-ray
The XRT onboard Swift (Burrows et al., 2005) performed follow-up observations in response to our target-of-opportunity (ToO) requests, with exposures ranging from 1.3 to 3.6 ks depending on the GRB, and all data were collected in Photon Counting (PC) mode. X-ray data, including light curves, calibrated event files, and spectra, were retrieved from the public Swift/XRT GRB Catalogue hosted by the UK Swift Science Data Centre [^5] https://www.swift.ac.uk/ and processed using the standard XRT pipeline as described in Evans et al. (2007, 2009). All light curves and spectra were generated using the automated tools provided by the XRT team.
XRT detected X-ray afterglows for all seven events, with uncatalogued sources coincident with or close to the GOTO optical transient locations in each case. The corresponding 0.3–10 keV light curves (in counts s-1) for the seven GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B, together with comparison GRBs, and spectra (in counts s-1 keV-1) for our sample are discussed later in Section 6.1.
4.2 UV/Optical/NIR Photometric Observations
Follow-up afterglow observations in the UV/optical/NIR, including data from both the GCN Circulars and this work, are compiled in Table LABEL:tab:phot_data. The table provides details on the observing facilities, instruments used, measured magnitudes, and other relevant parameters. Figure 6 shows the extinction-corrected UV/optical/NIR afterglow light curves and the wavelength coverage of all filters used (passbands in nm).
While GOTO primarily contributed discovery optical observations for the GRBs in our sample, its relatively small aperture and lack of multiple band observations limit its utility for extended follow-up. Details of the GOTO observations and initial detection magnitudes have been presented in Section 3.2. Therefore, GOTO observations are not included in this section. Instead, this section focuses on subsequent optical afterglow follow-up observations obtained with a range of facilities situated around the globe.
4.2.1 Swift/UVOT
In addition to XRT, Swift simultaneously observes with its UV-Optical Telescope (UVOT; Roming et al. 2005). We obtained the resulting data from the UK Swift Science Data Centre [^6] https://www.swift.ac.uk/index.php (UKSSDC) and used uvotproduct v2.9777As part of HEASoft v6.32 (Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), 2014). to measure the photometry of the afterglow. We used a 5″ radius circular aperture centred at the positions noted in Section 3.2 and a detection threshold of $3\sigma$. The measured magnitudes were converted from the UVOT photometric system to AB using the standard UVOT zeropoints (Breeveld et al., 2011). The afterglow was detected in at least one epoch for five of the seven sources in our sample. The exceptions are GRB 240112A, where the UVOT FoV didn’t cover the source position, and GRB 240225B, for which no UVOT data were obtained. The UVOT follow-up observations are listed in Table LABEL:tab:phot_data, and the corresponding light curves are presented in Figure 6.
4.2.2 LT
The IO:O Imager at the robotic 2m Liverpool Telescope (LT; Steele et al., 2004) located at the international Observatorio del Roque de los Muchachos, La Palma, was triggered for follow-up of GRBs 240122A, 240225B, and 240910A. For GRBs 240122A and 240910A, one epoch was obtained in each of rizriz bands. For GRB 240225B, 1 epoch is in grizgriz, and three later epochs were secured in $r$-band. All LT data were pre-reduced for bias, dark, and flat-field corrections using the facility pipeline. The photometry was then extracted with the photometry-sans-frustration pipeline (psf; Nicholl et al., 2023), making use of its built-in template subtraction. The observations are summarised in Table LABEL:tab:phot_data, and the corresponding light curves are shown in Figure 6.
Figure 6: UV/optical/NIR afterglow photometry for GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B, combining our measurements with values compiled from GCN Circulars (see Table LABEL:tab:phot_data). Left ordinates show flux density (Jy) and right ordinates show AB magnitude; corrected for foreground extinction. Downward arrows indicate non-detections ($3\sigma$ upper limits). To reduce crowding, per-band vertical magnitude offsets are applied as noted in each legend; an offset of $\Delta m$ corresponds to a multiplicative factor of $10^{-0.4\Delta m}$ on the flux axis. Time is measured relative to the trigger ($T{-}T_{0}$): we adopt the MAXI $T_{0}$ for GRBs 240122A and 240225B and the Fermi $T_{0}$ for the remaining five bursts. The bottom-right panel summarises the approximate passband coverage (nm) of all UV/optical/NIR filters used.
4.2.3 NOT
The Nordic Optical Telescope (NOT; Djupvik & Andersen 2010) is a 2.56 m telescope located at the Observatorio del Roque de los Muchachos in La Palma (Canary Islands, Spain). The NOT routinely performs ToO observations of GRB and FXT afterglows. The NOT observed and detected the counterparts of GRBs 240225B, 240619A, 240910A, 240916A, and 241228B using the Alhambra Faint Object Spectrograph and Camera (ALFOSC) optical imager. The reduction of the NOT data follows standard procedures, including bias and flat-field correction. The photometric calibration was computed against the Pan-STARRS catalogue. Details of these measurements are provided in Table LABEL:tab:phot_data, with their temporal evolution illustrated in Figure 6.
4.2.4 VLT/Xshoter acquisition camera
The X-shooter spectrograph (Vernet et al., 2011) mounted on Unit Telescope 3 (Melipal) of the Very Large Telescope (VLT) at Cerro Paranal Observatory was triggered for follow-up of a subset of GRBs in our sample.
Prior to the spectroscopic observations (see Section 4.3.1), acquisition images were obtained with the acquisition and guiding (A&G) camera in the Sloan $g^{\prime}$, $r^{\prime}$, and $z^{\prime}$ bands, depending on the target. These images were used to verify the target acquisition and also provide valuable photometric information on the afterglow. The raw frames were reduced using a custom pipeline based on ccdproc (Craig et al., 2017), including bias subtraction and flat-field correction.
Astrometric calibration was applied using Astrometry.net (Lang et al., 2010), and the images were aligned and stacked where appropriate. Aperture photometry was performed using photutils (Bradley et al., 2024), and the zero-point was calibrated against Pan-STARRS DR2 field stars. A log of the observations is given in Table LABEL:tab:phot_data, and the resulting light curves are presented in Figure 6.
4.2.5 BOOTES
The Burst Observer and Optical Transient Exploring System (BOOTES) [^8] https://bootesnetwork.com/ observed GRB 240122A using the 60 cm robotic telescope at the BOOTES-2/TELMA station in La Mayora, Málaga, Spain. Observations began on 2024 January 22 at 19:44:56 UT, approximately 9.3 hours after the trigger. A series of 60 s exposures were obtained with a clear filter. The afterglow was faint and remained undetected in both individual frames and the stacked image. A $3\sigma$ upper limit was derived, as listed in Table LABEL:tab:phot_data.
4.2.6 OSN
The follow-up observations of GRB 240122A are also performed with the 1.5 m telescope at the Sierra Nevada Observatory (OSN, Granada, southern Spain) [^9] http://www.osn.iaa.es/, targeting the burst position starting on 2024 January 22 at 23:45:57 UT (13.3hr post-trigger). The afterglow was clearly detected during the first night, prompting continued monitoring over the following two nights (January 23 and 24). Observations across all three epochs were performed in the Johnson–Cousins $B$, $V$, $R$, and $I$ bands, with exposure times of 90 s and 150 s. The afterglow remained clearly visible in the stacked images. Photometric measurements were obtained via aperture photometry using standard procedures in the IRAF software package (Tody, 1986), following bias subtraction and flat-field correction. Magnitudes were calibrated against nearby reference stars in the field, listed in the SDSS catalogue, using the transformation equations from Lupton et al. (2005) [^10] http://www.sdss.org/dr4/algorithms/sdssUBVRITransform.htm. The resulting magnitudes are reported in Table LABEL:tab:phot_data, while Figure 6 displays the corresponding light curves.
4.2.7 1.5m AZT-33IK Mondy
The AZT-33IK 1.5-meter telescope at the Sayan Solar Observatory (ISTP SB RAS), located near the village Mondy in Buryatia, was triggered for the follow-up observations of GRB 240225B on 2024-02-27 at 13:01:23 UT. A series of 30 images with an individual exposure of 120 seconds was obtained using the Andor NEO CMOS photometer attached to the telescope. The observations were carried out using the Johnson R-filter. Aperture photometry of the stacked image from the entire series yielded a magnitude for the optical afterglow of $R=19.72\pm 0.07$ (AB). We continued observations with AZT-33IK until the 2024-03-17 epoch.
We processed all our observations using the APEX pipelines (Devyatkin et al., 2010; Kouprianov, 2012; Pankov et al., 2022). The process involved image calibration (dark frame subtraction, flat-fielding, cosmic-ray removal), image quality control, image stacking, and source extraction. The apex_forced_phot pipeline was utilised for the forced photometry of the GRB 240225B afterglow on difference images. Image subtraction with the method described in Tomaney & Crotts 1996, was performed by apex_subtract pipeline, using PanSTARRS-DR1 survey images as a reference obtained from HIPS2FITS service [^11] https://alasky.cds.unistra.fr/hips-image-services/hips2fits (Boch et al., 2020). This step ensured that the underlying host galaxy, presented in the Legacy Survey DR9 with a magnitude of $r\sim 24.2$ and a photometric redshift of $z\sim 0.9$, did not affect the afterglow measurement.
We note an LS DR10 source at the coordinates RA = 08h33m26s.06 and Dec = +270432′′.9 (8′′ West of the afterglow position) with a magnitude of $r\sim 22.4$ that may affect the photometry in the images with poor seeing. The apparent magnitudes were initially calibrated against three nearby USNO-B1.0 stars (identifiers are 1171-0194062, 1171-0194079, and 1171-0194031) in the Vega system and then converted to the AB system using standard Vega-to-AB magnitude conversion [^12] https://www.astronomy.ohio-state.edu/martini.10/usefuldata.html. The complete record of these observations is compiled in Table LABEL:tab:phot_data, and their light curves are plotted in Figure 6.
4.2.8 LBT
We obtained late-time $r^{\prime}z^{\prime}$ imaging of GRB 240910A with the Large Binocular Cameras (LBCs; Giallongo et al., 2008) mounted on the Large Binocular Telescope (LBT) on Mt. Graham, Arizona, USA (Program ID: IT-2024B-023). LBC imaging data were reduced using the dedicated data reduction pipeline (Fontana et al., 2014). Aperture photometry was performed via IRAF tools and calibrated against SDSS field stars. Observation logs are reported in Table LABEL:tab:phot_data, with the associated light-curve behaviour shown in Figure 6.
4.2.9 HMT
The Half-Meter Telescope (HMT) is a 50 cm wide-field telescope, located at Nanshan Station of Xinjiang Astronomical Observatory, Chinese Academy of Sciences. HMT conducted two observations during the night of 2024-02-26 (UT), 18.40 and 24.87 hr since the GRB 240225B trigger, respectively. Based on the standard data processing of IRAF and aperture photometry, the measured brightness in R-band is listed in Table LABEL:tab:phot_data, and plotted in Figure 6.
4.2.10 TRT
The Thai Robotic Telescope (TRT) is an automated telescope network comprising four 70 cm CDK700 Telescopes equipped with Andor CCD cameras, distributed in the United States (SRO), Chile (CTO), Australia (SBO), and China (GAO). The telescope located at SBO started observations at 8.7 hr after the GRB 241228B trigger and obtained $4\times 300$ s frames in the $R$-band. The Johnson-Cousin filters were calibrated with the converted magnitude from the Sloan system [^13] https://www.sdss4.org/dr12/algorithms/sdssUBVRITransform/#Lupton. The aperture photometry is calibrated with the Pan-STARRS Data Release 2 (Chambers et al., 2016; Flewelling, 2018) and listed in Table LABEL:tab:phot_data.
4.2.11 Altay
The Altay Telescopes are located at the Altay Observatory, Xinjiang, China, as part of the Altay Astronomical Time-domain Project (also known as JinShan Project). This project consists of four 50 cm telescopes with a FOV of $1.7\degree\times 1.7\degree$, which are named from 50A to 50B, two 100 cm telescopes with a FOV of $1.4\degree\times 1.4\degree$, which are named 100A and 100B, and one 100 cm telescope with a FOV of $14^{\prime}\times 14^{\prime}$, which is named 100C. In the early commissioning stage of the project, we triggered the GRB 240910A and GRB 240916A with the 100 cm telescopes using the Sloan $r$-filter.
The obtained images were processed with the standard IRAF procedures, including bias and dark subtraction, flat correction, and image combination. After the astrometric calibration by Astrometry.net (Lang et al., 2010), the apparent photometric data were calibrated with the Pan-STARRS Data Release 2 (Chambers et al., 2016; Flewelling, 2018). A comprehensive summary of the results is given in Table LABEL:tab:phot_data, and light curves are shown in Figure 6.
4.2.12 1.6m Mephisto
The 1.6m Multi-channel Photometric Survey Telescope (Mephisto) is a wide-field multichannel telescope (Yuan et al., 2020). It is located at Lijiang Observatory of Yunnan Astronomical Observatories, Chinese Academy of Sciences, and is operated by the South-Western Institute for Astronomy Research, Yunnan University. Equipped with three-channel CCD cameras (blue $uv$, yellow $gr$, and red $iz$ channels), Mephisto can perform simultaneous observations in $ugi_{m}$ or $vrz_{m}$ optical bands at a particular moment. The wavelength coverage of the $u_{m},v_{m},g_{m},r_{m},i_{m}$, and $z_{m}$ filters is 320–365, 365–405, 480–580, 580–680, 775–900, and 900–1050 nm with central wavelengths at 345, 385, 529, 628, 835, and 944 nm, respectively (see, e.g., Yang et al., 2024; Cheng et al., 2025). Presently, the facility is in an advanced stage of commissioning.
Mephisto was triggered to observe GRB 241228B on 2024-12-28 (15:38:16) UT and continued until 2024-12-31. Multiple frames with an exposure time of 300 s were obtained at different epochs during the follow-up. The pre-processing of raw frames was performed using a specialised pipeline developed for the Mephisto observational data (Fang et al., in prep.). To obtain the instrumental magnitudes of the GRB, Point Spread Function photometry was performed on the stacked images. The corrected Gaia XP low-resolution spectra (Huang et al., 2024) were utilised for the photometric calibration. Considering that the Mephisto bands are not fully covered by the corrected Gaia XP spectra (336–1020 nm), it was extrapolated partially in the $u$ and $z$ Mephisto bands. Each band’s synthetic magnitude in the AB system was calculated by convolving the spectra with the transmission efficiency. The median of the magnitude offset between the instrumental and synthetic magnitudes of the non-variable stars in the field was used to finally calibrate the Mephisto photometric measurements (for details see Chen et al., 2024; Zou et al., 2025). The overall uncertainties in the photometric calibration were constrained to be within 0.03, 0.01, and 0.005 mag in the $u_{m}$, $v_{m}$, and $griz_{m}$ bands, respectively. The detailed dataset is compiled in Table LABEL:tab:phot_data, and the corresponding light curves are shown in Figure 6.
4.3 Spectroscopic Observations
For the GRBs in our sample, we acquired spectra using the X-shooter instrument mounted on the VLT (Vernet et al., 2011), and OSIRIS (Optical System for Imaging and lowIntermediate-Resolution Integrated Spectroscopy) on the Gran Telescopio Canarias (GTC; Cepa 1998), except for GRB 241002B, for which we did not get any spectroscopic observation due to scheduling constraints. A complete summary of the VLT/X-shooter and GTC/OSIRIS spectroscopic configurations, exposure details, and observing conditions is provided in Table 3, whereas spectra are shown in Figures (A1)–(A7).
Table 3: Log of spectroscopic afterglow observations for the GRBs analysed in this work. Redshift values are included here for completeness; detailed measurement methods behind estimating these values are described in Section 6.3.
GRB Date-Obs T-T0 Telescope Instrument Exp. Time Slit width Airmass Seeing Redshift
(hours) (′′) (′′)
240122A 2024-01-22 22:56:23 UT 12.47 GTC OSIRIS $3\times 900$ s 1.0 1.50 1.70 $3.1634\pm 0.0003$
240122A 2024-01-23 02:25:01 UT 15.95 VLT X-shooter $4\times 1200$ s 1.0a-0.9b 1.00-1.01 0.45-0.49 $3.1634\pm 0.0003$
240225B 2024-02-29 01:45:36 UT 77.52 VLT X-shooter $4\times 1200$ s 1.0a-0.9b 1.61-1.65 0.66-0.70 $0.9462\pm 0.0002$
240619A 2024-07-02 23:20:52 UT 331.68 VLT X-shooter $4\times 600$ s 1.0a-0.9b 1.90-2.00 0.94-0.96 $0.3960\pm 0.0001$
240910A 2024-09-12 03:08:33 UT 47.13 GTC OSIRIS $3\times 1200$ s 1.0 1.14 0.80 $1.4605\pm 0.0007$
240916A 2024-09-17 00:08:31 UT 22.76 VLT X-shooter $4\times 600$ s 1.0a-0.9b 1.74-1.99 1.07-1.09 $2.6100\pm 0.0002$
241228B 2024-12-29 06:00:45 UT 25.68 VLT X-shooter $4\times 1200$ s 1.0a-0.9b 1.17-1.18 0.53-0.54 $2.6745\pm 0.0004$
NOTE: aUVB arm, bVIS and NIR arms.
4.3.1 VLT/X-shooter spectrograph
Spectroscopic observations of GRBs 240122A, 240225B, 240619A, 240916A, and 241228B were performed as part of the “Stargate” GRB program at ESO, using the X-shooter spectrograph (Vernet et al., 2011), installed on the ESO VLT UT3 at Cerro Paranal, Chile. X-shooter simultaneously covers the ultraviolet-blue (UVB; 300–560 nm), visible (VIS; 550–1020 nm), and near-infrared (NIR; 1020–2100 nm) wavelength ranges, with resolving powers of $\lambda/\Delta\lambda=5400$, $8900$, and $5600$, respectively. Observations were carried out in the ABBA nodding mode along the slit to enable effective subtraction of the sky emission, especially in the NIR. In addition, a K-band blocking filter was employed to reduce thermal background contamination in the NIR. The data were reduced in STARE mode using calibration files from the night and the standard ESO X-shooter pipeline (Goldoni et al., 2006; Modigliani et al., 2010), which performs bias and dark correction, flat-fielding, wavelength calibration via arc lamps, and flux calibration based on standard star observations. Following the method described in Selsing et al. (2019), individual reduced exposures were directly co-added for the UVB and VIS arms, while A–B nod pairs were pair-subtracted prior to combination for the NIR arm. All reported wavelengths are given as observed in vacuum and corrected for the barycentric motion of the Earth.
The X-shooter observations of GRB 240122A began on 2024 January 23 at 02:25:01 UT ($T_{0}+15.95$ hr) under excellent seeing conditions of $0.49^{\prime\prime}$ (Saccardi et al., 2024). GRB 240225B was observed on 2024 February 29, starting at 01:45:36 UT ($T_{0}+3.23$ d), with a seeing of 0.66′′ (Schneider et al., 2024). We observed GRB 240619A on 2024 July 02 starting at 23:20:52 UT ($T_{0}+13.82$ d); our observations targeted the catalogued galaxy PSO J162.3946+17.2828 in spatial coincidence with the optical afterglow and were obtained with a seeing of $0.94^{\prime\prime}$ (Cotter et al., 2024). Spectroscopic observations of GRB 240916A were conducted on 2024 September 16, beginning at 00:08:31 UT ($T_{0}+22.76$ hr), under a seeing of $1.07^{\prime\prime}$ (Pieterse et al., 2024). Finally, X-shooter observations of GRB 241228B were carried out on 2024 December 29, starting at 06:00:45 UT ($T_{0}+1.07$ d), with a seeing of $0.54^{\prime\prime}$ (An et al., 2024).
In all, for GRBs 240122A, 240225B, 240619A, 240916A, 241228B observed with VLT/X-shooter, the start times of the observations spanned across the sample from $T_{0}+15.95$ hr to $T_{0}+13.82$ d, seeing ranged from $0.49^{\prime\prime}$ to $1.07^{\prime\prime}$, and each target was obtained in four exposures per arm with per-exposure times of either 600 s or 1200 s; for GRB 240619A the slit also encompassed a nearby second galaxy. The full spectroscopic observing log, along with observing conditions, is listed in Table 3. The final column lists the precise redshifts estimated for these GRBs, derived from the analysis of their afterglow spectra. These values were determined through the identification of absorption and/or emission features associated with the host galaxies. The full methodology, including line identification, fitting procedures, and associated uncertainties, is described in detail in Section 6.3.
4.3.2 GTC/OSIRIS
Spectroscopic observations of GRBs 240122A and 240910A were performed using the OSIRIS mounted on the 10.4 m GTC (Cepa, 1998) at the Observatorio del Roque de los Muchachos (ORM), La Palma (see full spectroscopic log in Table 3). The observations were obtained as part of GTC programs GTCMULTIPLE2J-23B and GTCMULTIPLE4G-24B (PI: J. F. Agüí Fernández). For both targets, the R1000B grism was used in long-slit spectroscopy mode (LSS) with a slit width of $1.0^{\prime\prime}$ and binning of $2\times 2$ pixels, providing a resolving power of $R\sim 600$ and a wavelength coverage of 3650–7800 Å.
The data were acquired in a sequence of three individual exposures, nodding along the slit to cancel the effect of possible artefacts or defects and provide a clean, final reduced product. The OSIRIS spectrum of GRB 240122A was began on 2024 January 22 at 22:56:23 UT ($T_{0}+12.47$ hr), approximately 3.5 hr prior to the VLT/X-shooter observations (Thoene et al., 2024), , with $3\times 900$ s exposures. The GTC/OSIRIS spectrum of GRB 240910A was started on 2024 September 12 at 03:08:33 UT ($T_{0}+47.13$ hr), using $3\times 1200$ s exposures, under good observing conditions (airmass $\sim 1.14$ and seeing of $0.8^{\prime\prime}$, de Ugarte Postigo et al. 2024).
Data reduction was carried out using a combination of IRAF-based tasks and custom Python scripts developed for OSIRIS, which included bias subtraction, flat-fielding, wavelength calibration using arc lamps, and flux calibration using spectrophotometric standards. Accurate one-dimensional spectra were extracted using optimal extraction techniques and corrected for instrumental response across the full wavelength range. Later on, Section 6.3 details the methodology, including line identification, fitting, and uncertainty estimation.
4.4 Radio
[…]
4.4.1 ATCA
Following its optical localisation, GRB 240122A was observed with the ATCA under the PanRadio GRB programme C3542 (PI: Anderson) on 2024 January 24, 26, 28, and February 12. This program aims to perform high-cadence multi-frequency radio monitoring of a large sample of LGRBs in the southern hemisphere (Declinations <10<-10 deg) between minutes to years after the burst to explore the evolution and properties of their afterglows (Leung et al., in prep.; Anderson et al., in prep.). GRB 240122A was observed with a wide range of frequencies centred on 5.5, 9.0, 16.7, and 21.2 GHz, each with a 2048 MHz-wide band. We reduced the visibility data using standard routines in MIRIAD (Sault et al., 1995). We used a combination of manual and automatic RFI flagging before calibration, conducted with MIRIAD tasks uvflag and pgflag, respectively. We used PKS 1934-63 to determine the bandpass response and to calibrate the flux density scale for all frequency bands. We used PKS 0607-157 to calibrate the time-variable complex gains for all epochs and frequency bands. After calibration, we inverted the visibilities using a robust weighting of 0.5 and then used the CLEAN algorithm (Clark, 1980) to the target source field using standard MIRIAD tasks INVERT, CLEAN, and RESTOR to obtain the final images. For each observation, we measure the flux density of a detected source by fitting a point-source model to the restored image using the Miriad task imfit and report a non-detection using the rms sensitivity obtained from the residual image. The $1\sigma$ errors reported are purely statistical, as the systematic errors are expected to be much smaller ($\lesssim 5$ per cent; e.g., Reynolds, 1994; Tingay et al., 2003). We detected the radio counterpart at 9 GHz on both 2024 January 24 and 28 at a position consistent with the GOTO optical counterpart (Kumar et al., 2024b). For all other frequencies, we estimated $3\sigma$ upper limits (see Table A2).
4.4.2 VLA
We observed GRB 240910A with the VLA 3.1 (2024 September 13, Giarratana et al. 2024a), 9.1 (2024 September 19), 21.3 (2024 October 1) and 46.2 (2024 October 26) days post-burst (Project code: SF171028) at the central frequencies of 6 (C band), 10 (X band) and 15 GHz (Ku band), with a bandwidth of 4, 4 and 6 GHz, respectively. The VLA source J0125-0005 was used as a phase calibrator. The distance between the target and the phase calibrator was about 2.7. Each observation included scans on the flux and bandpass calibrator 3C48. The data were calibrated using the custom casa pipeline (Version 6.5.4; McMullin et al., 2007) and visually inspected for possible radio frequency interference. The final images were produced with the tclean task in casa (Version 6.5.4) using a Briggs weighting scheme (robust =0.5=0.5). Results from the campaign are reported in Table A2. The GRB is detected at all frequencies during the first two epochs, and at 6 GHz it is also detected in the third epoch. The maximum flux densities, measured in the first epoch, are $37\pm 10, 11$$14\pm 9$, and $86\pm 10$ uJy at 6, 10, and 15 GHz, respectively. For each detection, the flux density was measured by fitting a Gaussian to the cleaned image using the imview task in casa. The final flux density error was estimated as the squared sum of the root mean square (RMS) and a typical 5% accuracy for the amplitude scale calibration. Upper limits are reported with a $5\sigma$ confidence level.
Table 4: Results from fitting the prompt emission spectra of MAXI/GSC and Fermi/GBM detected GRBs. The models used are those preferred based on the Akaike Information Criterion (AIC, Akaike 1974). In order to maximise the number of measured peak energies, the cutoff power-law (CPL) and Band (Band et al., 1993) models were chosen when AIC was agnostic between them and the PL model. Hardness ratios (HR) are calculated using the 50–300 keV and 10–50 keV bands. HR values for MAXI/GSC bursts are from extrapolating the 2–20 keV fits to these energies.
GRB $z$ $T_{\rm 90}$ Model Fluence [2 – 20 keV] $\alpha$ $\beta$ $E_{\rm p}$ PG-Statistic DoF HR
(MAXI/GSC) (s) ($10^{-7}$ erg cm-2) (keV)
240122A $3.163$ $\approx 36$ PL $1.26^{+0.47}_{-0.94}$ $1.9\pm 0.4$ - - $1.32^{+1.30}_{-0.65}$
240225B $0.946$ $\approx 21$ PL $10.7^{+0.63}_{-1.26}$ $1.9\pm 0.1$ - - $1.32^{+0.25}_{-0.21}$
GRB $z$ $T_{\rm 90}$ Model Fluence [10 – 1000 keV] $\alpha$ $\beta$ $E_{\rm p}$ PG-Statistic DoF HR
(Fermi/GBM) (s) ($10^{-5}$ erg cm-2) (keV)
240619A $0.396$ $36.13\pm 0.59$ Band $1.29_{-0.08}^{+0.05}$ $1.36^{+0.14}_{-0.17}$ $1.65^{+0.09}_{-0.04}$ $149.5_{-143.5}^{+423.7}$ 241.14 436 $2.28_{-0.07}^{+0.06}$
240910A $1.460$ $272.39\pm 2.61$ Band $2.16_{-0.24}^{+0.18}$ $1.25^{+0.08}_{-0.05}$ $2.81^{+2.83}_{-0.44}$ $113.3_{-22.7}^{+27.5}$ 271.38 324 $1.81_{-0.07}^{+0.06}$
240916A $2.610$ $32.00\pm 0.81$ Band $2.51_{-0.15}^{+0.11}$ $1.11^{+0.04}_{-0.06}$ $2.16^{+0.32}_{-0.14}$ $665.6_{-186.1}^{+231.0}$ 231.02 325 $4.26_{-0.11}^{+0.08}$
241002B - $64.26\pm 4.38$ PL $0.43_{-0.07}^{+0.06}$ $1.97^{+0.06}_{-0.08}$ - - 219.85 326 $1.18_{-0.11}^{+0.09}$
241228B $2.674$ $19.46\pm 0.36$ CPL $4.19_{-0.13}^{+0.10}$ $0.81^{+0.03}_{-0.02}$ - $350.6_{-22.9}^{+17.9}$ 310.74 325 $5.72_{-0.07}^{+0.06}$
5 Prompt Emission Analysis and Properties
The prompt gamma-ray emission encodes the immediate output of the central engine and provides key diagnostics of the physical conditions in the relativistic outflow. For the seven GRBs in our sample, we analysed MAXI/GSC and Fermi/GBM data to measure basic spectral and temporal properties, including the photon indices, peak energies ($E_{\rm p}$), isotropic-equivalent gamma-ray energies ($E_{\rm iso}$), and $T_{\rm 90}$; analyses using other high-energy instruments discussed in Section 3.1 are beyond the scope of this work. These quantities are critical for placing the bursts in the broader GRB population, identifying any outliers, and examining potential links between the prompt emission and the optical afterglows recovered by GOTO. In particular, we aim to investigate whether the unusually hard spectra and high $E_{\rm p}$ values observed in several cases are connected to the optical detectability of these poorly localised events.
5.1 Prompt Emission Analysis
Here, we describe the methods used to extract and analyse the prompt emission properties for each GRB, using data from the relevant high-energy instruments. The analysis is divided into two parts: the MAXI/GSC events and the Fermi/GBM events.
5.1.1 GRBs 240122A and 240225B with MAXI/GSC
We analysed the prompt emission of GRBs 240122A and 240225B with archival data of MAXI/GSC using High Energy Astrophysics Software (HEASOFT [^14] https://heasarc.gsfc.nasa.gov/docs/software/heasoft/). X-ray events of gsc_med type are processed with mxproduct. Because light curves and spectra produced by mxproduct are not suitable for short and variable transients like GRBs, we performed an additional process to extract light curves with a 1-second time resolution and applied effective area correction. Note that the process is identical to the process used for the MAXI GRB catalogue.
Figure A8 shows the GSC light curves of GRBs 240122A and 240225B in the 2–20, 2–4, 4–10, and 10–20 keV energy bands. We estimated the $T_{\rm 90}$ duration in the 2–20 keV band to be $\approx 36$ s and $\approx 21$ s for GRBs 240122A and 240225B, respectively. For GRB 240225B, the duration was calculated using data from a single MAXI/GSC scan. This value differs from those obtained by other instruments because the scan began at 20:15:30 UT, about 200 s later than the trigger times reported by the others.
Then, we extract a spectrum of the $T_{\rm 90}$ interval and correct for variations in the effective area. The spectra of GRBs 240122A and 240225B are fit with a single power-law model, and the photon indices are found to be $1.9\pm 0.4$ and $1.9\pm 0.1$ ($1\sigma$ error), respectively. The energy flux in 2–20 keV was $3.5_{-2.6}^{+1.3}\times$10-9 erg cm-2 s-1 and $5.1_{-0.6}^{+0.3}\times$10-8 erg cm-2 s-1. Figure A9 shows the spectrum with the best-fit model. The results of each fit are shown in Table 4.
5.1.2 GRBs 240619A, 240910A, 240916A, 241002B, and 241228B with Fermi/GBM
We analysed the prompt emission of GRBs 240619A, 240910A, 240916A, 241002B, and 241228B using the Fermi/GBM data available from the High Energy Astrophysics Science Archive Research Center (HEASARC [^15] https://heasarc.gsfc.nasa.gov) archive (von Kienlin et al., 2020; Gruber et al., 2014; von Kienlin et al., 2014; Bhat et al., 2016). Using the HEASOFT and the Fermi Gamma-ray Data Tools (Goldstein et al., 2023), we took the time-tagged event (TTE) data in the $T_{\rm 90}$ interval for each burst to use as the source. We used the data from the brightest NaI detectors and the corresponding BGO detectors. These were NaI 0, NaI 1, NaI 2 and BGO 0 for GRB 240619A; NaI 8, NaI 11, and BGO 1 for GRB 260910A; NaI 3, NaI 4, and BGO 0 for GRB 240916A; NaI 10, NaI 11, and BGO 1 for GRB 241002B; and NaI 6, NaI 7, and BGO 1 for GRB 241228B. The background was modelled in the standard way, using a polynomial function fit to the CSPEC data for each burst. The order of the polynomial was increased until the reduced fit statistic was <1.15<1.15, resulting in the models fitting each GRB suitably well. Once the background model was obtained, we interpolated it across the source interval and exported this as the background to be used for spectral analysis. Additionally, we extracted the necessary response files.
We performed our analysis of the spectra using PyXspec (Gordon & Arnaud, 2021) using three models of varying complexity: a simple power-law (PL), which measures a photon index, $\alpha$; a power-law with a high-energy exponential cutoff (CPL), which measures $\alpha$ and the spectral peak energy, $E_{\rm p}$; and the Band model (Band et al., 1993), which measures two photon indices $\alpha$ and $\beta$, smoothly connected at a characteristic break energy, $E_{c}$. This break energy is converted to a peak energy using $E_{p}=E_{c}(2-\alpha)$. We used the PG-statistic in our analysis, which is appropriate for Poisson data with a Gaussian background. The results of each fit are shown in Table 4. Figures A10 and A11 present the Fermi/GBM observations of GRBs 240619A, 240910A, 240916A, 241002B, and 241228B, showing the light curves and the corresponding fitted spectra, respectively.
5.2 Prompt Emission Properties
We plot the 4/5 Fermi bursts where it was possible to measure both $E_{\rm p}$ and $E_{iso}$ (see Table 4) on the Amati plane (Amati et al., 2002; Amati, 2006) in Figure 7. GRBs 240916A and 241228B appear to show unusually high redshift-corrected peak energies ($E_{p,i}$) relative to their $E_{iso}$ measurements, and are inconsistent with the Amati relation at the $3\sigma$ level. Both bursts lie at the high end of the $E_{iso}$ distribution, indicating intrinsically powerful GRBs. In addition to this, the high energy photon index ($\beta$) measured for GRB 240619A is less than 2, indicating that the true peak of the spectrum is at an even higher energy. This may make 240619A an outlier of the Amati relation too. Only GRB 240910A appears to be typical in terms of its measured prompt properties.
We also plot all five Fermi and two MAXI GRBs on an HR – $T_{\rm 90}$ diagram (c.f. Kouveliotou et al., 1993), see Figure 8. It is immediately apparent that the Fermi GRBs are unusually spectrally hard. Only GRB 241002B (best fit with a power-law model with a photon index of $\alpha=1.97^{+0.08}_{-0.06}$) sits within the main ‘cloud’ of LGRBs. The others are either longer in $T_{\rm 90}$ (GRB 240910A), or driven to high HR by their abnormally hard (for collapsar GRBs) photon indices (GRBs 240619A, 241228B) and/or high $E_{\rm p}$ (GRB 240916A). The MAXI GRBs also sit at the high end of the GBM HR distribution, but these values are obtained by extrapolating the 2–20 keV spectral fits to 300 keV, and so should be considered upper limits because we have no constraints on any breaks in the spectrum.
Figure 7: Fermi GRBs in our sample plotted in the Amati plane (Amati et al., 2002; Amati, 2006), showing the relationship between the intrinsic (redshift corrected) peak energy ($E_{p,i}$) and the isotropic-equivalent gamma-ray energy release ($E_{iso}$). Lines show the best-fit correlation (solid) and their $3\sigma$ bounds (dashed) for the long (blue) and short (grey) GRB populations. Correlation fits and comparison data are from Minaev & Pozanenko (2020).
Figure 8: Hardness ratio vs $T_{\rm 90}$ for the five Fermi/GBM bursts in our sample. The hardness ratio is the 50 – 300 keV fluence over the 10 – 50 keV fluence. Comparison data is taken from von Kienlin et al. (2020).
Figure 9: Distribution of $T_{\rm 90}$ durations for a sample of GRBs from the BAT catalogue, divided into long (blue) and short (grey) populations. The dashed vertical lines mark the durations of GOTO GRBs, labelled with their respective names and $T_{\rm 90}$ values. All GOTO GRBs fall in the long-duration class and cover a wide range of durations.
Given that the GRBs in our sample were selected based on the recovery of their optical afterglows by a relatively shallow telescope like GOTO, it’s perhaps not surprising that they’re outliers with respect to the wider sample of GRBs. GRBs 240916A and 241228B both appear to be straightforward cases of intrinsically powerful GRBs, as evidenced by their higher than average $E_{iso}$. However, their measured $E_{p,i}$ is high relative to the Amati relation even when accounting for the large $E_{iso}$.
In contrast, GRB 240619A appears to be a much more energetically typical GRB, but at a low redshift ($z=0.3960$). The higher hardness ratio in this case may simply be the result of less redshifting than the majority of the detected population, resulting in a harder portion of the synchrotron spectrum falling in the 10–300 keV bandpass than usual. GRB 240910A also appears to be more typical energetically. In this case, a longer central engine duration ($T_{90}=272.39\pm 2.61$ s) may be responsible for producing the bright afterglow.
All four of the above GRBs (240619A, 240910A, 240916A and 241228B) have measured low-energy photon indices of $\alpha<1.5$. This indicates that the low-energy synchrotron break was likely in the bandpass, causing $\alpha$ to be an unusually hard blend of two portions of the synchrotron spectrum (see, e.g., Ravasio et al., 2018, 2019). The positions of these breaks are functions of the underlying physical parameters of the jet, and unusual parameter values may be responsible for the abnormally high hardness ratio. An alternative explanation is that the unusually hard prompt spectra in GRB 240916A and GRB 241228B may reflect jet–stellar-envelope interactions, supported by structured-jet and radiative-transfer simulations of LGRBs (Lazzati & Begelman, 2005; Lazzati et al., 2013; Lundman et al., 2013) and even low-luminosity jets (Aloy et al., 2018).
5.3 $T_{90}$ comparison
The $T_{\rm 90}$ duration is a key parameter for classifying GRBs into long and short populations, with a conventional threshold at $T_{90}=2$ s. Figure 9 shows the distribution of $T_{\rm 90}$ values from the BAT GRB catalogue [^16] https://swift.gsfc.nasa.gov/results/batgrbcat/index_tables.html, plotted on a logarithmic scale and separated into long (blue) and short (grey) classes. The histogram illustrates the well-known bimodality in the GRB population, with the majority of events falling in the long-duration category.
Overlaid on this distribution are the durations of the seven GOTO-detected GRBs in our sample, marked with vertical dashed lines and annotated with their names and $T_{\rm 90}$ values. All seven events lie securely within the long-duration class, with durations ranging from $\sim 20$ s to $\sim 272$ s. GRB 241228B ($19.5$ s) and GRB 240225B ($21$ s) sit at the lower end of the LGRB population, while GRBs 240122A, 240619A, 240916A, and 241002B (326432-64 s) are closer to the peak of the LGRB distribution. GRB 240910A ($272$ s) lies toward the higher end, placing it among the longest events in the Swift sample.
Table 5: Summary of Swift/XRT afterglow spectral and temporal properties for our GRB sample. Exp. denotes the spectral extraction time interval. $N_{\rm H,Gal}$ is fixed to the Galactic line-of-sight value, while $N_{\rm H,int}$ is the intrinsic absorption component derived from spectral fitting. Fluxes are unabsorbed values in the 0.3–10.0 keV band. Negative $\alpha$ values indicate apparent rising trends likely caused by low-count statistics.
[…]
($10^{3}$ s) ($10^{22}$ cm-2) ($10^{22}$ cm-2) ($10^{-12}$ erg cm-2 s-1) (s)
240122A 29.5 $2.50^{+0.70}_{-2.10}$ $1.88^{+0.31}_{-0.24}$ 0.105 $\sim$0.0 $6.54^{+1.41}_{-1.16}$ 926.5 3.163
240225B 461.7 $1.09^{+0.16}_{-0.13}$ $2.19^{+0.72}_{-0.59}$ 0.040 $0.29^{+0.49}_{-0.29}$ $12.6^{+6.5}_{-3.2}$ 289.7 0.946
240619A 174.0 $0.94^{+3.69}_{-0.05}$ $1.61^{+0.45}_{-0.71}$ 0.028 $0.038^{+0.034}_{-0.038}$ $1.23^{+0.54}_{-0.36}$ 1978.0 0.396
240910A 128.3 $1.10^{+0.40}_{-0.30}$ $1.75^{+2.22}_{-0.61}$ 0.027 $0.006^{+0.005}_{-0.006}$ $0.69^{+0.49}_{-0.27}$ 1983.0 1.460
240916A 61.0 $1.29^{+0.34}_{-0.23}$ $2.90^{+0.80}_{-0.60}$ 0.140 $2.17^{+2.63}_{-1.92}$ $4.34^{+3.92}_{-1.52}$ 1643.0 2.610
241002B 134.5 $1.10^{+0.40}_{-0.30}$ $1.82^{+1.06}_{-0.70}$ 0.030 $0.096^{+0.251}_{-0.096}$ $1.34^{+0.68}_{-0.37}$ 2702.0
241228B 40.0 $1.28^{+4.49}_{-0.03}$ $1.44^{+0.46}_{-0.32}$ 0.003 <2.37<2.37 $2.81^{+1.65}_{-0.69}$ 1809.0 2.674
6 Afterglow Analysis and Properties
In this section, we analyse the afterglow properties of the GOTO-discovered GRBs across X-ray, UV and optical (photometric and spectroscopic), and radio wavelengths. We characterise their temporal and spectral behaviour, and place the results in the context of the broader GRB afterglow population.
6.1 X-ray
Each Swift/XRT-detected source confirmed its association with the GOTO optical GRB counterpart on the basis of spatial coincidence and, in most cases, temporal fading. For time-domain analysis, we utilised the automated Swift/XRT light curve fits from the UKSSDC pipeline (Evans et al. (2007, 2009)). The X-ray light curves were adequately described by single power-law decays (see Figure 10 for the light curves in counts per second).
Spectra for both source and background regions, along with corresponding ancillary and response files extracted in PC mode from the Swift/XRT repository, were grouped to a minimum of one count per bin using grppha task. Spectral fitting was performed in XSPEC using Cash statistics (Cash, 1979), appropriate for low-count Poisson-distributed data. Each spectrum was modelled with a simple absorbed powerlaw, tbabs*ztbabs(zpowerlw) (Wilms et al., 2000), accounting for both Galactic foreground and intrinsic absorption in the host galaxy. Galactic column density ($N_{\rm H,g}$) for each GRB was fixed at a value obtained from Swift Galactic $N_{\rm H}$ tool [^17] https://www.swift.ac.uk/analysis/nhtot/ (Willingale et al., 2013), while the intrinsic host galactic absorption ($N_{\rm H,intr}$) and photon spectral index ($\Gamma$) were left free to vary. All fits were performed in the 0.3–10.0 keV energy range, and the errors reported here correspond to 90% confidence intervals. Figure A12 presents the 0.3–10.0 keV XRT count rate spectra for all seven GRBs in our sample, overlaid (red solid line) with their respective best-fit absorbed powerlaw models.
Table 5 summarises the temporal and spectral properties derived from Swift/XRT observations of our GRB sample. Redshifts used for spectral fitting are reported in later sections. Photon indices lie in the range $\Gamma\simeq 1.4-2.9$, consistent with typical afterglow spectra. The intrinsic absorption shows substantial variation: GRBs 240122A and 240910A are consistent with negligible additional absorption beyond the Galactic foreground, whereas others, most notably GRB 240916A, require higher column densities of the order of $10^{22}$ cm-2. The inferred unabsorbed 0.3–10 keV fluxes span nearly two orders of magnitude, from $\sim 7\times 10^{-13}$ to $\sim 1.3\times 10^{-11}$ erg cm-2 s-1, with the highest values observed for GRBs 240225B and 240122A, despite their relatively short spectral extraction intervals. Temporal decay slopes cluster around $\alpha\approx 1.0-1.3$ for most afterglows, though three cases (GRBs 240619A, 240910A, and 241228B) show formally flat or rising indices; in each case, the uncertainties are large and the behaviour is consistent with constant flux within errors. At such late phases, exposure times were modest (typically 1–3 ks), naturally limiting the statistical precision of the spectral fits.
Figure 10: Swift/XRT 0.3–10 keV light curves. Archival LGRBs (grey) and SGRBs (blue dashed) are plotted as background. The seven GOTO GRBs are shown with distinct colours and styles.
Figure 11: $R$- and $rr^{\prime}-$band afterglow light curves for the seven GRBs in our sample compared against archival GRB afterglows compiled by Dainotti et al. (2024), known as the "Kann plot".
6.1.1 Swift/XRT light curves comparison
To investigate the temporal behaviour of GRB afterglows and compare our GOTO-discovered GRBs against known populations, we compiled a comprehensive sample of Swift/XRT light curves, combining both archival GRBs and the GOTO sample (Figure 10). Light curves were obtained from the UK Swift Science Data Centre (UKSSDC) [^18] https://www.swift.ac.uk/xrt_products/bulk.php bulk access portal, using a custom notebook provided by the UKSSDC team to automate downloads. In our visualisation, archival LGRBs are shown in blue solid lines, SGRBs in grey dashed lines, and the seven GOTO-discovered GRBs (240122A, 240225B, 240619A, 240910A, 240916A, 241002B, 241228B) are overplotted with distinct markers and colours.
As seen in Figure 10, the GOTO GRBs lie well within the canonical LGRB distribution, with count rates and temporal slopes consistent with typical long-burst afterglows. Their X-ray light curves track the faint to intermediate range of the LGRB population, showing no evidence for extreme behaviour. GRBs 240225B, 240910A, 240916A and 241002B in particular seems to follow smooth declines, while others (e.g., GRB 241228B) are represented only by a few points, underscoring the sparse nature of the coverage. This sparseness arises not from the afterglows themselves but from the fact that these bursts were not initially triggered by Swift/BAT or XRT, but instead by wide-field, poorly localised instruments such as Fermi/GBM and MAXI/GSC.
6.2 UV/Optical/NIR
UV/Optical/NIR afterglow light curves for all seven GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B in our sample are shown in Figure 6 (see Table LABEL:tab:phot_data for complete photometric observations), plotted in both flux density and apparent magnitude space. For visual clarity, magnitudes in different filters have been offset vertically where indicated in the legends. All magnitudes are reported in the AB system and have been corrected for Galactic extinction using the E(BV)E(B-V) values listed in Table 1, based on the recalibrated dust maps of Schlafly & Finkbeiner (2011).
Among the GRBs in our sample, GRB 241228B shows the highest cadence multi-colour coverage, with early bright detections in GOTO $L$-band and a well-sampled decline. GRBs 240122A and 240225B each have moderate multi-band coverage, while events such as GRBs 241002B and 240916A have sparser datasets but still provide key temporal constraints. The sparse and uneven temporal coverage of the light curves prevents us from robustly constraining decay slopes, break times, or colour evolution. In most cases, only a few photometric points are available per GRB, which precludes detailed afterglow fitting. Nevertheless, these datasets enable comparison with the extended sample of GRB afterglow light curves discussed in the following subsection and provide the basis for the afterglow modelling presented in Section 7.
6.2.1 Optical light curves comparison - Kann plot
To compare the optical behaviour of the GOTO-detected GRBs in our sample with previously well-observed GRB afterglows, we plotted their light curves alongside a reference sample from Dainotti et al. 2024 (Figure 11). Three GRBs (240122A, 240225B, 241228B) have well-sampled $R$-band data, while the remaining four (240619A, 240910A, 240916A, 241002B) are well-observed in the $rr^{\prime}$ band (see Figure 6). In each case, the same filter band from the reference sample is used to minimise colour offsets, and all magnitudes are corrected for Galactic extinction. The reference light curves span a wide range of brightness and decay behaviours, and are plotted in grey for comparison.
The left panel of Figure 11 shows the $R$-band events (240122A, 240225B, 241228B) in blue, red, and lime, while the right panel shows the $r/r^{\prime}$-band events (240619A, 240910A, 240916A, 241002B) in purple, magenta, maroon, and navy. The afterglow light curves comparison in $R$- and $rr^{\prime}$ bands demonstrates that the afterglows of GOTO-detected GRBs lie within the overall distribution of known GRB afterglows in both brightness and decay behaviour. The $R$-band events are consistent with the median properties of the sample, whereas the $rr^{\prime}-$band events span a broader range in brightness and decline rates.
To place our events quantitatively within the broader population, we interpolated the extinction–corrected light curves at fixed epochs, using only those with data coverage near the corresponding epochs of our GRBs. In $R$ band at $t=0.79$ d (close to the first data point for GRB 240225B), the comparison sample spans 14.1414.1423.9323.93 mag (median $20.45$ mag). Our GRBs fall within this range: GRB 240225B is relatively bright ($18.79$ mag), while GRBs 241228B and 240122A are near the population median ($20.49$ and $20.98$ mag, respectively). In $rr^{\prime}$ band at $t=1.64$ d (a phase where both GRBs 240910A and 240916A have measurements), the sample spans 14.2014.2025.5025.50 mag (median $21.13$ mag). GRB 240916A is somewhat brighter than average ($19.90$ mag), while GRB 240910A lies close to the median ($20.80$ mag) but shows an unusually steep decline, fading by $\Delta m\approx 4.7$ mag over $\Delta t\approx 14.7$ d ($\sim$0.32 mag d-1). By contrast, GRB 241002B appears relatively faint even at early times. Overall, these comparisons confirm that the GOTO afterglows occupy the central brightness distribution of known GRB afterglows while also sampling the diversity of decline rates and brightness within the population.
6.3 Spectroscopic Analysis – Redshift Estimation
When a GRB explodes, the resulting afterglow light passes through both the interstellar medium of its host galaxy and any intervening material along the line of sight, imprinting a series of absorption features onto the spectrum. In our X-shooter spectra, only the highest redshift absorption system is identified and assigned as the redshift of the GRB since no higher redshift intervening material is physically possible. While additional foreground absorption systems may be present, a detailed analysis and characterisation of these intervening absorbers is beyond the scope of this paper and will be addressed in future work. The redshifts are estimated by identifying common absorption lines in GRB afterglows using the line lists of Fynbo et al. (2009) and Christensen et al. (2011), and/or emission lines from their host galaxies. The redshift and its associated uncertainty are then derived by fitting Voigt profiles Krogager (2018) to the absorption features, prioritizing low-ionization, unsaturated, and unblended transitions, and Gaussian profiles to the emission lines.
Figure 12: Histogram of redshift distribution for long (blue) and short (grey) GRBs from the sample of Dainotti et al. (2024). Overlaid are the redshifts of six GRBs with optical afterglows discovered by GOTO in 2024, marked with vertical dashed lines, labelled by GRB name within the plot, and redshift values above. This visual comparison highlights the diversity in redshift of GOTO-discovered GRBs and demonstrates their placement within the broader GRB population.
[…]
6.3.1 VLT/X-shooter
Here, we summarise the results from our VLT/X-shooter spectra of GRBs 240122A, 240225B, 240619A, 240916A and 241228B:
GRB 240122A
[…]
GRB 240225B
In the case of GRB 240225B, a continuum is detected over the entire wavelength range (from 3300 to 20400 Å) and the following strong absorption features are identified: Al iii, Cr ii, Fe ii, Mn ii, Mg ii, Mg i, and Ca ii at a common redshift of $z=0.9462\pm 0.0002$. At the same redshift, three emission lines ([O ii] $\lambda\lambda 3727,\,3730$ and [O iii] $\lambda 5008$) are identified from the host galaxy. The complete list of identified lines is provided in Table A4. One intervening system is identified at $z=0.7056$.
GRB 240619A
[…]
GRB 240916A
[…]
GRB 241228B
[…]
6.3.2 GTC/OSIRIS
Here, we summarise the results from our GTC spectra of GRBs 240122A and 240910A:
GRB 240122A
[…]
GRB 240910A
In the case of GRB 240910A, the afterglow continuum is clearly detected across the full spectral range, and the spectrum reveals a rich set of absorption features. Prominent lines include Si ii, C iv, Fe ii, Al ii, Al iii, Cr ii, Mn i, Mn ii, Ni ii, Mg ii, and Mg i, along with several fine-structure transitions such as Fe ii and Ni ii. All lines are consistent with a common redshift of $z=1.4605\pm 0.0007$. The reduced spectrum is presented in Figure A5, and a complete list of identified lines is provided in Table A6.
In summary, our spectroscopic follow-up of seven GRBs using VLT/X-shooter and GTC/OSIRIS reveals a wide range of redshifts ($z\sim 0.40$ to $z\sim 3.16$), with absorption and emission features tracing both the GRB host environments and the intervening interstellar medium. High-quality afterglow spectra enable precise redshift measurements and identification of various ionic species, including fine-structure transitions. These results provide critical context for understanding the physical conditions in GRB host galaxies and lay the foundation for future studies of metallicity, dust content, and kinematics in GRB environments.
6.3.3 Redshift Comparison
To show the redshift distribution of GRBs in our sample, we compare them against the broader GRB population presented in the comprehensive compilation by Dainotti et al. (2024). Figure 12 shows a histogram of GRBs with measured redshifts from that sample, classified into long and short categories. The majority of GRBs in the Dainotti et al. (2024) sample are LGRBs, with a redshift distribution peaking around $z\sim 0.5-2$, consistent with the star formation history of the universe. SGRBs appear more frequently at lower redshifts, consistent with their likely origin from compact object mergers with longer delay times.
Overlaid on this distribution are the measured redshifts of the GOTO GRBs, shown as vertical dashed lines with annotations above the axis. Our sample spans from $z=0.40$ (GRB 240619A) to $z=3.16$ (GRB 240122A). It is worth noting that the optical selection imposed by GOTO inherently limits detections to $z\lesssim 5$, since at higher redshifts the Lyman forest progressively enters and then highly absorbs flux across the GOTO $L$-band (400–700 nm). GRBs 240225B ($z=0.95$) and 240910A ($z=1.46$) fall near the central peak of the LGRB distribution, while GRB 240916A ($z=2.61$) and GRB 241228B ($z=2.67$) occupy the higher-redshift tail together with GRB 240122A at $z=3.16$. This spread highlights that the GOTO sample encompasses both low- and high-redshift GRBs, demonstrating the survey’s capability to probe the wide observed redshift range of the LGRB population.
From a physical perspective, the low-redshift events, such as GRB 240619A, are particularly valuable for detailed host-galaxy and supernova connection studies, where high signal-to-noise follow-up is achievable. Conversely, the higher-redshift events (e.g., GRBs 240916A, 241228B, and 240122A) provide critical leverage for probing star-forming environments in the early universe and for constraining the role of GRBs as tracers of cosmic star formation beyond $z>2$.
Table 6: Parameter estimation priors and marginalised posteriors for the GOTO-discovered GRBs using the afterglowpy TopHat model. Posteriors are medians with $16-84\%$ credible intervals.
GRB 240122A; GOTO24eu
$\theta_{v}$ (rad) $\log_{10}(E_{0})$ (erg) $\theta_{c}$ (rad) $\log_{10}(n_{0})$ (cm-3) $p$ $\log_{10}\epsilon_{e}$ $\log_{10}\epsilon_{B}$ $\xi_{N}$ $d_{L}$ (Mpc)
Priors U(0.0; 0.5) U(49; 57) U(0.0; 0.5) -0.379 2.119 -1.246 -4.290 1.0 27708.87
Posteriors $2.15_{-0.10}^{+0.09}\times 10^{-2}$ $54.97_{-0.02}^{+0.02}$ $2.39_{-0.12}^{+0.12}\times 10^{-2}$ -0.379 2.119 -1.246 -4.290 1.0 27708.87
GRB 240225B; GOTO24tz
$\theta_{v}$ (rad) $\log_{10}(E_{0})$ (erg) $\theta_{c}$ (rad) $\log_{10}(n_{0})$ (cm-3) $p$ $\log_{10}\epsilon_{e}$ $\log_{10}\epsilon_{B}$ $\xi_{N}$ $d_{L}$ (Mpc)
Priors U(0.0; 0.5) U(49; 57) U(0.0; 0.5) U(-5.0; 3.0) 2.119 -1.246 -4.290 1.0 6342.55
Posteriors $9.43_{-0.41}^{+0.38}\times 10^{-2}$ $54.27_{-0.04}^{+0.04}$ $7.23_{-0.31}^{+0.31}\times 10^{-2}$ $1.87_{-0.11}^{+0.09}$ 2.119 -1.246 -4.290 1.0 6342.55
GRB 240619A; GOTO24cvn
$\theta_{v}$ (rad) $\log_{10}(E_{0})$ (erg) $\theta_{c}$ (rad) $\log_{10}(n_{0})$ (cm-3) $p$ $\log_{10}\epsilon_{e}$ $\log_{10}\epsilon_{B}$ $\xi_{N}$ $d_{L}$ (Mpc)
Priors U(0.0; 0.5) U(49; 57) U(0.0; 0.5) -0.379 2.119 U(-5.0; 0.0) U(-5.0; 0.0) 1.0 2217.13
Posteriors $0.41_{-0.07}^{+0.04}$ $52.70_{-0.35}^{+0.42}$ $0.45_{-0.07}^{+0.04}$ -0.379 2.119 $0.85_{-0.39}^{+0.32}$ $2.18_{-0.15}^{+0.12}$ 1.0 2217.13
GRB 240910A; GOTO24fvl
$\theta_{v}$ (rad) $\log_{10}(E_{0})$ (erg) $\theta_{c}$ (rad) $\log_{10}(n_{0})$ (cm-3) $p$ $\log_{10}\epsilon_{e}$ $\log_{10}\epsilon_{B}$ $\xi_{N}$ $d_{L}$ (Mpc)
Priors U(0.0; 0.5) U(49; 57) U(0.0; 0.5) -0.379 2.119 -1.246 -4.290 0.180 10826.91
Posteriors $4.25_{-0.12}^{+0.07}\times 10^{-2}$ $54.77_{-0.01}^{+0.01}$ $4.07_{-0.12}^{+0.08}\times 10^{-2}$ -0.379 2.119 -1.246 -4.290 1.0 10826.91
GRB 240916A; GOTO24fzn
$\theta_{v}$ (rad) $\log_{10}(E_{0})$ (erg) $\theta_{c}$ (rad) $\log_{10}(n_{0})$ (cm-3) $p$ $\log_{10}\epsilon_{e}$ $\log_{10}\epsilon_{B}$ $\xi_{N}$ $d_{L}$ (Mpc)
Priors U(0.0; 0.5) U(49; 57) U(0.0; 0.5) -0.379 2.119 U(-5; 0) U(-5; 0) 0.180 22000.43
Posteriors $2.29^{+0.71}_{-0.78}\times 10^{-2}$ $54.40^{+1.03}_{-0.63}$ $4.73^{+1.16}_{-1.26}\times 10^{-2}$ -0.379 2.119 $1.34^{+0.59}_{-0.95}$ $1.47^{+0.91}_{-0.73}$ 1.0 22000.43
GRB 241228B; GOTO24jmz
$\theta_{v}$ (rad) $\log_{10}(E_{0})$ (erg) $\theta_{c}$ (rad) $\log_{10}(n_{0})$ (cm-3) $p$ $\log_{10}\epsilon_{e}$ $\log_{10}\epsilon_{B}$ $\xi_{N}$ $d_{L}$ (Mpc)
Priors U(0.0; 0.5) U(49; 57) U(0.0; 0.5) -0.379 2.119 -1.246 -4.290 0.180 22653.17
Posteriors $4.86^{+0.14}_{-0.11}\times 10^{-3}$ $55.429^{+0.004}_{-0.004}$ $10.56^{+0.05}_{-0.07}\times 10^{-3}$ -0.379 2.119 -1.246 -4.290 0.180 22653.17
NOTE: $\theta_{v}$ – viewing angle; $E_{0}$ – isotropic-equivalent kinetic energy; $\theta_{c}$ – half-opening angle of jet core; $n_{0}$ – density of the surrounding ISM; $p$ – electron energy
distribution power law index; $\epsilon_{e}$ – fraction of energy that goes into electrons; $\epsilon_{B}$ – fraction of energy that goes into the magnetic field; $\xi_{N}$ – fraction of
shock-accelerated electrons; $d_{L}$ – luminosity distance.
GRB 240122A
GRB 240225B
GRB 240619A
GRB 240910A
GRB 240916A
GRB 241228B
Figure 14: Multi-wavelength afterglow light curves for six GRBs modelled in this work: GRBs 240122A, 240225B, 240619A, 240910A, 240916A, and 241228B. Each panel shows the observed data points (UV/optical/NIR, X-ray, and radio, where available), overlaid with the best-fit afterglow model using a TopHat (uniform) jet scenario (see text for details). Photometric data are compiled from our observations and published GCN Circulars. Fluxes are rescaled for clarity as indicated in the legends.
GRB 240122A
GRB 240225B
GRB 240619A
GRB 240910A
GRB 240916A
GRB 241228B
Figure 15: Posterior distributions for the TopHat-jet afterglow model parameters inferred for six GRBs analysed in this work: GRBs 240122A, 240225B, 240619A, 240910A, 240916A, and 241228B. Each panel shows the marginalised 1D and 2D posterior distributions from the dynesty nested-sampling run (via Bilby), with contours representing $68\%$ and $95\%$ credible regions. Inferred parameters include the observer angle, core angle, isotropic-equivalent energy, and microphysical quantities (see text for details).
6.4 Radio
The radio light curves of GRBs 240122A, 240619A, 240910A, and 240916A are shown in Figure 13, plotted in flux density (left) and luminosity (right). For context, we compare these with the historical population of GRBs compiled at 8–10 GHz. The GOTO GRBs lie within the locus of LGRBs, showing flux densities and luminosities consistent with this population. None displays the systematically fainter or more rapidly fading behaviour typical of SGRBs. Within the sample, GRB 240122A is among the brightest radio afterglows, while GRB 240910A and GRB 240916A fall at the lower end of the distribution, illustrating the intrinsic spread in LGRB radio emission. Despite this variation, their temporal evolution remains broadly consistent with expectations for LGRB afterglows, reinforcing the conclusion that the radio properties of the GOTO events trace the same underlying population.
7 Afterglow modelling
For the afterglow light curve modelling, we used the afterglowpy module (version 0.8.0; Ryan et al. 2020). This Python-based tool utilises the single-shell approximation (van Eerten et al., 2010; van Eerten, 2018) to model GRB afterglow light curves by accounting for the effects of complex jet structures and an off-axis observer position. We modelled multi-band light curves for each of the GOTO-discovered GRBs presented in this paper. For our modelling, we assumed the simplest TopHat profile of the jet structure and fixed some parameters where required (see Table 6).
All afterglow analyses were conducted using dynesty nested sampling within the Bilby framework (version 2.4.0; Ashton et al. 2019). The dataset mostly consisted of relatively sparse data in optical bands, complemented by X-ray data from Swift/XRT. To ensure robustness and minimise bias, our priors were set to be broad, and the prior probabilities for most parameters were modelled using uniform distributions (see Table 6). We used a Gaussian likelihood and the dynesty nested sampler with 1500 live points and a stopping tolerance of $\Delta\log\mathcal{Z}=0.1$.
When modelling GRBs with sparsely sampled afterglow data, we adopt a flexible strategy in which a subset of parameters is fixed to literature-informed medians that are representative of LGRBs drawn from previous population studies (Aksulu et al., 2022; Chrimes et al., 2022) to assess consistency with the typical long-GRB population. These works analysed large samples of LGRBs and reported values for key microphysical parameters: $\epsilon_{e}$ (Aksulu et al., 2022), and $\epsilon_{B}$, $n_{0}$, and $p$ (Chrimes et al., 2022). A value for $\log_{10}\epsilon_{e}$ is not reported in Chrimes et al. (2022), from which we take most other fixed parameters owing to its larger sample. We therefore adopt $\epsilon_{e}$ from Aksulu et al. (2022). We do not take all parameters from Aksulu et al. (2022) because its GRB sample is smaller; combining their $\epsilon_{e}$ with the broader Chrimes et al. (2022) set maximises coverage. Our adopted value is slightly below the peak $\epsilon_{e}\sim 0.13-0.15$ inferred from radio peaks by Beniamini & van der Horst (2017), but it lies within the $\epsilon_{e}\sim 0.01-0.16$ range for a homogeneous medium reported by Duncan et al. (2023), who used radio peaks together with constraints from the prompt gamma-ray emission efficiency.
For GRBs where full sampling over all parameters led to unconstrained, multimodal, or non-convergent posteriors, we fixed one or more of these quantities to their literature-based mean values. This approach allows stable and interpretable modelling when the data cannot independently constrain all parameters. Fixing select values based on well-motivated priors reduces degeneracies, avoids overfitting, and maintains physical plausibility in the resulting fits.
Results overview. We modelled six GOTO-discovered GRBs (excluding GRB 241002B; no redshift) with the TopHat (uniform) jet model in afterglowpy using dynesty nested sampling via Bilby. Posterior summaries (medians with $16-84\%$ credible intervals) are listed in Table 6; multi-band light curves and posterior corner plots are shown in Figures 14 and 15, respectively. Unless stated otherwise, we adopt $\xi_{N}=1$ (fraction of electrons accelerated) as our baseline; for sparsely constrained events, we fix a subset of microphysical parameters ($p$, $n_{0}$, and, where noted, $\epsilon_{B}$; see Table 6) to population-informed values to suppress degeneracies. With this setup, the events are well described by narrow, near–on-axis geometries; expected covariances (e.g., $E_{0}-\theta_{c}$) are present but posteriors are unimodal and not prior-bound.
For completeness, we provide short remarks on a subset of GRBs in our sample for which extra clarification is useful. These notes highlight only the key features or caveats, while the overall methodology and global results are presented above.
GRB 240122A: A three-parameter TopHat fit ($\theta_{v}$, $\theta_{c}$, $\log_{10}E_{0}$) with microphysics fixed as explained above reproduces the broadband evolution (Table 6; Figures 1415). Radio points lie slightly above the model, while late-time X-ray points are slightly below; since synchrotron self-absorption would further suppress early-time radio, it cannot explain the positive radio residuals—more plausible are a reverse shock, mild energy injection, a density bump, or calibration offsets.
GRB 240910A: A three-parameter TopHat fit ($\theta_{v}$, $\theta_{c}$, $\log_{10}E_{0}$) with the other microphysics held fixed reproduces the optical and X-ray light curves with a single parameter set (Table 6; Figures 1415). The model shows a modest, systematic overprediction in the radio bands. This behaviour is consistent with fixed microphysics — at fixed $\epsilon_{e}$, adopting $\xi_{N}=1$ raises $F_{\nu,\max}$ and lowers $\nu_{m}$, which naturally boosts the radio while leaving higher-frequency bands close to the data. For completeness, we also explored fits in which additional microphysical parameters were allowed to vary, including $\xi_{N}$; these trials slightly reduced the radio residuals but degraded the X-ray agreement and produced broader, strongly correlated posteriors. For uniformity across the sample, we therefore retain the three-parameter fit and interpret the residual radio offsets as secondary systematics related to normalisation and propagation effects (e.g., synchrotron self-absorption; Sari et al. 1998; Granot & Sari 2002) and interstellar scintillation (Goodman 1997; Frail et al. 1997); possible host free–free absorption is also plausible (see, e.g., Weiler et al. 2002).
GRB 240916A. A five-parameter TopHat (uniform) jet fit ($\theta_{v}$, $\theta_{c}$, $\log_{10}E_{0}$, $\log_{10}\epsilon_{e}$, $\log_{10}\epsilon_{B}$) yields a clean broadband match with unimodal posteriors. The geometry is nearly on-axis with small angles. The microphysics favour a comparatively lower $\epsilon_{e}$ and a moderately higher $\epsilon_{B}$ within typical afterglow ranges. Magnetised internal–shock models for the prompt phase indicate that magnetisation can alter radiative efficiency and shift the characteristic synchrotron/IC spectral peaks (e.g. Mimica & Aloy, 2012). However, the $\epsilon_{B}$ inferred here is the downstream magnetic energy fraction of the afterglow forward shock and is not directly comparable to the prompt–phase shell magnetisation; any putative link to $E_{p,i}$ is therefore model-dependent and not required by our data.
GRB 241002B: No secure redshift is available. We sampled $z$ with a broad prior, but the redshift posterior remained unconstrained; redshift-dependent quantities ($d_{L}$, $E_{0}$, and rest-frame times) track the priors, and the angles are only weakly informed. We therefore do not report parameter estimates and exclude this burst from population-level comparisons.
GRB 241228B. We model the afterglow with a three–parameter TopHat jet, sampling $(\theta_{v},\ \theta_{c},\ \log_{10}E_{0})$ while holding the microphysics fixed ($p$, $n_{0}$, $\epsilon_{e}$, $\epsilon_{B}$, $\xi_{N}$). This minimal configuration reproduces the optical and X–ray evolution at early–to–intermediate epochs (Table 6; Figures 1415). Small, band–dependent residuals appear around the X–ray band and the bluest optical filters, consistent with a cooling break lying close to the optical and/or modest host galaxy extinction; these offsets remain at a low level under reasonable microphysical choices. Because the microphysics are fixed, the fit can trade flux normalisation against geometry, and the posterior favours an effectively on–axis view with a very narrow core. We therefore regard the recovered $\theta_{c}$ as a model–dependent lower bound — allowing, for example, $\epsilon_{B}$ to vary with a broad log–uniform prior would broaden the $\theta_{c}$ posterior and plausibly shift its median upward, at the cost of a higher $E_{0}$.
At late times, the model systematically underpredicts the flux across bands, indicating additional physics not captured by a single forward–shock component. Plausible explanations include mild, sustained energy injection (refreshed shocks), interaction with a local density enhancement, or an additional outflow component (e.g., a wider jet or cocoon). These effects can maintain the late–time emission above the single–jet prediction without disrupting the early–time agreement. Given the limited leverage to discriminate among scenarios, we retain the three–parameter TopHat fit for the main analysis and note the late–time excess as a likely secondary component.
Beaming factor and jet energetics. For each GRB in our sample (see Table 6) we compute the beaming factor and the beaming-corrected jet energy using the relations $f_{b}=1-\cos\theta_{c},\qquad E_{\rm jet}=f_{b}\,E_{0}.$
Per-burst summary (medians with 16–84% uncertainties):
240122A:  $f_{b}=(2.86^{+0.31}_{-0.28})\times 10^{-4}$,  $E_{\rm jet}=(2.66^{+0.20}_{-0.19})\times 10^{51}\,\mathrm{erg}$.
240225B:  $f_{b}=(2.61^{+0.22}_{-0.22})\times 10^{-3}$,  $E_{\rm jet}=(4.83^{+0.06}_{-0.05})\times 10^{51}\,\mathrm{erg}$.
240619A:  $f_{b}=(9.79^{+1.76}_{-2.73})\times 10^{-2}$,  $E_{\rm jet}=(4.80^{+6.51}_{-2.51})\times 10^{51}\,\mathrm{erg}$.
240910A:  $f_{b}=(8.29^{+0.33}_{-0.50})\times 10^{-4}$,  $E_{\rm jet}=(4.82^{+0.15}_{-0.21})\times 10^{51}\,\mathrm{erg}$.
240916A:  $f_{b}=(1.12^{+0.62}_{-0.52})\times 10^{-3}$,  $E_{\rm jet}=(2.89^{+14.20}_{-1.93})\times 10^{51}\,\mathrm{erg}$.
241228B:  $f_{b}=(5.58^{+0.05}_{-0.08})\times 10^{-5}$,  $E_{\rm jet}=(1.50^{+0.01}_{-0.01})\times 10^{51}\,\mathrm{erg}$.
Overall, the parameters inferred from our TopHat (uniform) jet fits — specifically the observer angle ($\theta_{v}$), the jet half-opening angle ($\theta_{j}\equiv\theta_{c}$ in this model), and the isotropic-equivalent kinetic energy ($E_{0}$) — are consistent with the ranges reported in previous studies of long-duration GRB afterglows. In all cases, we recover $\theta_{v}$ and $\theta_{j}$ of order a few degrees, in line with the narrow jet geometries commonly found in broadband afterglow fits (e.g. Frail et al. (2001); Bloom et al. (2003); Friedman & Bloom (2005); Racusin et al. (2009)). We use $E_{0}$ to estimate the beaming-corrected jet kinetic energy, yielding values that cluster around $10^{51}-10^{52}$ erg, consistent with the canonical long-GRB energy scale (e.g. Friedman & Bloom (2005)). This agreement in both angular geometry and energetics supports the robustness of our TopHat-jet modelling and places these events firmly within the established population of classical long-duration GRBs.
8 Summary and Conclusion
Over the past few years, the Gravitational-wave Optical Transient Observer (GOTO) has become instrumental in the search for, and rapid follow-up of, optical counterparts to poorly localised transients such as GRBs and GW events. Since achieving first light in June 2017, GOTO has steadily progressed from its prototype (GOTO-4) phase into a fully operational dual-site facility (GOTO-36).
During the prototype era (2017–2020), the GOTO-4 system responded to 77 Fermi/GBM and 29 Swift/BAT triggers, securing its first optical afterglow detection with GRB 171205A. Following its expansion to GOTO-36, GOTO attempted follow-up observations of more than 257 Fermi, 43 Swift, 28 EP, and 7 GECAM triggers up to 31 December 2024. Whereas, to date, GOTO has issued nearly 80 GCN circulars and yielded $\approx 28$ confirmed afterglow detections, ranging from rapid identifications such as GRB 230818A within $4.43$ min of the trigger, to wide-field discoveries of poorly localised events including GRB 230911A and the SGRB 241105A. Collectively, these results highlight GOTO’s ability to respond on timescales as short as 36 seconds and to cover hundreds of square degrees in order to identify optical afterglows under challenging localisation conditions.
Within this broader context, the present study focuses on the first systematic sample of LGRB afterglows detected by GOTO, discovered during 2024. Our sample comprises seven LGRBs (GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B), including two MAXI/GSC events (GRBs 240122A and 240225B), localised to arcminute precision and detected serendipitously during survey operations, and five Fermi/GBM events (GRBs 240619A, 240910A, 240916A, 241002B, and 241228B), recovered through rapid, targeted tiling of degree-scale localisation regions. For all seven LGRBs, GOTO provided the earliest optical detections, with response times ranging from $\sim 19.3$ min to $9.4$ hr and with sky coverage exceeding $75\%$ of the $95\%$ probability regions for the GBM bursts.
Notably, GRB 241228B provides an illustrative example: its afterglow was identified on the 94.5% probability contour, outside the typical GBM $90\%$ localisation region. While the majority of GRB counterparts are recovered within the $90\%$ region, a small fraction are expected to lie beyond, making this a noteworthy case that highlights both the statistical nature of localisation regions and the importance of wide-field optical follow-up. These results highlight the adaptability and efficiency of the GOTO network in responding to both well-localised (MAXI/GSC) and more uncertain (Fermi/GBM) GRBs. Regardless of the size or shape of the localisation area, GOTO’s rapid tiling strategy enabled meaningful coverage and facilitated the identification of several optical afterglows within its fields.
These rapid identifications enabled immediate triggering of Swift/XRT and UVOT observations and coordinated multi-wavelength follow-up using facilities around the globe, underscoring the central role of optical discovery of poorly localised GRBs in constraining their properties. The follow-up campaign yielded detections in the X-ray, UV, optical, and radio bands for most of the events in our sample. Swift/XRT confirmed X-ray counterparts for all bursts. Optical photometry from multiple facilities provided light curves extending from minutes to days post-trigger, showing a broad range of brightnesses and decay rates. Spectroscopy for five events using the VLT/X-shooter (GRBs 240122A, 240225B, 240619A, 240916A, and 241228B) and for two events using GTC/OSIRIS (GRBs 240122A and 240910A) delivered precise redshifts spanning $z\approx 0.40-3.16$, along with absorption line diagnostics tracing both host galaxy interstellar media and, for the higher redshift bursts, intervening absorbers. Radio detections for four GRBs (240122A, 240619A, 240910A, and 240916A) utilising mainly ATCA and VLA confirmed long-lived synchrotron emission, most likely arising from forward shocks. Taken together, this multi-wavelength dataset has enabled robust classification and placed each burst in the broader context of the LGRB population.
Analysis of the prompt emission using MAXI/GSC and Fermi/GBM data revealed a spectrally hard sample, with four events yielding measurable $E_{\rm p}$ values (GRBs 240619A, 240910A, 240916A, and 241228B). Two bursts, GRBs 240916A and 241228B, stand out as $3\sigma$ outliers to the Amati relation, while others displayed unusually hard low-energy photon indices, pointing to diversity in jet microphysics and, in some cases, potentially high magnetisation. The measured $T_{\rm 90}$ durations, ranging from $\sim$20 s to over 270 s, confirm all seven events as LGRBs, encompassing both short-engine and long-engine members of the class.
Comparisons with the broader GRB population reinforce this conclusion. The GOTO-detected afterglows occupy the established luminosity–time phase space of LGRBs in both X-rays and optical, while their radio detections likewise follow the known locus of synchrotron afterglows. Their redshifts ($z\sim 0.40-3.16$) span both nearby and distant events. Taken together, these results demonstrate that the GOTO sample is representative of the wider population, probing afterglows across X-ray, optical, and radio wavelengths and capturing their diversity in temporal evolution and redshift. This highlights GOTO’s capability to deliver well-localised optical counterparts that integrate seamlessly with multi-wavelength studies of GRBs.
The contrast between typical afterglow behaviour and spectrally hard prompt emission in the GOTO sample likely reflects an observational bias: GRBs with higher $E_{\rm p}$ generally have larger $E_{\rm iso}$ and correspondingly brighter afterglows, making them easier to detect at optical wavelengths. While this tendency favours luminous events in poorly localised searches, it also provides a useful window into jet microphysics and central engine diversity. At the same time, it highlights the importance of wide-field optical facilities in complementing high-energy triggered samples and extending GRB studies across both nearby and high-redshift regimes.
We modelled the afterglows of six of seven GRBs in our sample (excluding GRB 241002B, which lacks a redshift). TopHat-jet parameters inferred here, observer angle ($\theta_{\mathrm{obs}}$), jet core angle ($\theta_{c}$), and isotropic-equivalent kinetic energy ($E_{0}$), are consistent with ranges typically found for LGRB afterglows. In all cases, we recover $\theta_{\mathrm{obs}}$ and $\theta_{c}$ of order a few degrees, in line with the narrow jet geometries commonly obtained from broadband afterglow fits. Using $E_{0}$ to estimate the true energy budget, the beaming-corrected jet kinetic energies cluster around $10^{51}-10^{52}$ erg, consistent with the canonical LGRB energy scale after correcting for beaming. Microphysical posteriors are broadly consistent with expectations for external-shock synchrotron emission. One event, GRB 241228B, shows a late-time flux excess relative to the best-fitting TopHat model, suggestive of an additional emission component or prolonged central-engine activity. Taken together, the geometry, energetics, and microphysics inferred from our uniform fits place these GOTO-discovered GRBs squarely within the established population of classical long-duration bursts while clarifying the levers that drive diversity in light-curve morphology. GOTO’s early discovery and dense optical cadence provide key leverage for constraining pre-break behaviour and for enabling robust, comparable modelling across events.
In all, the results presented in this study clearly demonstrate that GOTO’s wide-field, dual-site, fully robotic design, combined with adaptive trigger-specific strategies, is highly effective for bridging the gap between poorly localised high-energy triggers and the precise positions needed for multi-wavelength follow-up. The detections presented here highlight GOTO’s ability to recover GRB afterglows in a wide range of redshifts, localisation scales, and intrinsic properties, spanning both representative events and rare, energetically extreme outliers. In the emerging era of multi-messenger astronomy, GOTO’s demonstrated capability for rapid, deep optical searches makes it a critical asset for identifying and characterising counterparts to both gravitational wave events and gamma-ray bursts, thus advancing our understanding of the most energetic explosions in the Universe.
Acknowledgements
This work makes use of the Gravitational-wave Optical Transient Observer (GOTO) observations that led to the discovery of the optical afterglows presented in this study. GOTO project acknowledges support from the Science and Technology Facilities Council (STFC, grant numbers ST/T007184/1, ST/T003103/1, ST/T000406/1, ST/X001121/1 and ST/Z000165/1) and the GOTO consortium institutions; University of Warwick; Monash University; University of Sheffield; University of Leicester; Armagh Observatory & Planetarium; the National Astronomical Research Institute of Thailand (NARIT); University of Manchester; Instituto de Astrofísica de Canarias (IAC); University of Portsmouth; University of Turku.
This work is partly based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programmes 110.24CF and 114.27PZ (Co-PIs Tanvir, Malesani, Vergani).
It is also partly based on data obtained with the instrument OSIRIS, built by a Consortium led by the Instituto de Astrofísica de Canarias in collaboration with the Instituto de Astronomía of the Universidad Autónoma de México. OSIRIS was funded by GRANTECAN and the National Plan of Astronomy and Astrophysics of the Spanish Government.
It is also partly based on observations made with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias. The NOT data were obtained under program IDs 68-020, 69-023, and 70-507 (PIs Malesani, Fynbo, Xu). The National Radio Astronomy Observatory and Green Bank Observatory are facilities of the U.S. National Science Foundation operated under cooperative agreement by Associated Universities, Inc. Mephisto is developed at and operated by the South-Western Institute for Astronomy Research of Yunnan University (SWIFAR-YNU), funded by the “Yunnan University Development Plan for World-Class University" and “Yunnan University Development Plan for World-Class Astronomy Discipline". Based on observations made at the Observatorio de Sierra Nevada (OSN), operated by the Instituto de Astrofísica de Andalucía (IAA-CSIC).
AK and JRM are supported by the UK Science and Technology Facilities Council (STFC) Consolidated grant ST/V000853/1. BPG acknowledges support from STFC grant No. ST/Y002253/1. BPG and DO acknowledge support from The Leverhulme Trust grant No. RPG-2024-117. BS acknowledges the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE31-0011 (project PEGaSUS). DS is supported by the UK Science and Technology Facilities Council (STFC, grant numbers ST/T007184/1, ST/T003103/1, and ST/T000406/1). SG acknowledges support from the Istituto Nazionale di Astrofisica (INAF), project number: 1.05.24.07.04. BK acknowledges support from the “Special Project for High-End Foreign Experts," Xingdian Funding from Yunnan Province; the Yunnan Key Laboratory of Survey Science (grant 202449CE340002), and National Key Research and Development Program of China (grant 2024YFA1611603). BW is supported by the UKRI’s STFC studentship grant funding, project reference ST/X508871/1 RLCS and SM acknowledge support from The Leverhulme Trust grant RPG-2023-240. NH, POB and NRT acknowledge support from UKRI/STFC grant ST/W000857/1. JDL acknowledges support from a UK Research and Innovation Future Leaders Fellowship (MR/T020784/1). TLK acknowledges a Warwick Astrophysics prize post-doctoral fellowship made possible thanks to a generous philanthropic donation. BS and SDV acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE31-0011 (project PEGaSUS). MAA acknowledges support from grants PID2021-127495NB-I00 (funded by MCIN/AEI/10.13039/501100011033, EU), ASFAE/2022/026 (funded by MCIN, EU NextGenerationEU PRTR-C17.I1), and CIPROM/2022/13 (Generalitat Valenciana). DLC acknowledges support from the UK Science and Technology Facilities Council (STFC) grant number ST/X001121/1. AS acknowledges support by a postdoctoral fellowship from the CNES. AR acknowledges support from the INAF project Supporto Arizona & Italia. AMC and LC acknowledge support from the Irish Research Council Postgraduate Scholarship No. GOIPG/2022/1008. MEW and IW are supported by the UKRI Science and Technology Facilities Council (STFC) DBM is funded by the European Union (ERC, HEAVYMETAL, 101071865). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. NSP and ASP are grateful to the Russian Science Foundation (project no. 23-12-00220) for their partial support of the data reduction, analysis of data. EVK is grateful to the Ministry of Science and Higher Education of the Russian Federation for financial support of this work; the AZT-33IK telescope is part of the Center for Common Use "Angara".
Data Availability
Data can be shared upon request to the corresponding author.
Appendix
Table A1: Optical afterglow observations of GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B compiled within this work, along with those collected from the reported GCNs. All the tabulated magnitudes are in AB system.
$T-T_{0}$ (h) Instrument/Telescope Exp. Time Filter Mag Mag err Source
GRB 240122A (T0= 2460331.93615); GOTO24eu
0.728 GOTO-S $4\times 45$ s $L$ 17.581 0.037 This work
9.800 0.6m BOOTES-2 $14\times 60$ s clearclear >>17.81 This work
9.836 UVOT/Swift $481$ s $u$ >>20.4 Siegel et al. (2024)
12.318 OSIRIS/GTC $30$ s $r^{\prime}$ 20.431 0.115 This work
13.837 1.5m OSN $10\times 90$ s $V$ 20.73 0.18 This work
13.865 1.5m OSN $11\times 90$ s $B$ >>22.09 This work
13.892 1.5m OSN $10\times 90$ s II 20.73 0.08 This work
13.907 1.5m OSN $11\times 90$ s $R$ 20.94 0.09 This work
13.168 IO:O/LT $3\times 60$ s $r$ 20.396 0.207 This work
13.237 IO:O/LT $3\times 60$ s $i$ 20.489 0.200 This work
13.307 IO:O/LT $3\times 60$ s $z$ 20.461 0.255 This work
15.666 X-shooter/VLT $11\times 20$ s $r^{\prime}$ 20.652 0.033 This work
15.689 X-shooter/VLT $3\times 60$ s $r^{\prime}$ 20.610 0.046 This work
15.778 X-shooter/VLT $3\times 60$ s $g^{\prime}$ 21.277 0.026 This work
15.827 X-shooter/VLT $3\times 60$ s $z^{\prime}$ 20.494 0.029 This work
36.198 1.5m OSN $10\times 90$ s $B$ >>22.15 This work
36.225 1.5m OSN $10\times 90$ s $V$ 21.56 0.28 This work
36.253 1.5m OSN $10\times 90$ s $R$ 21.83 0.18 This work
36.280 1.5m OSN $10\times 90$ s II 22.05 0.15 This work
59.202 1.5m OSN $12\times 150$ s $B$ >>22.31 This work
59.247 1.5m OSN $12\times 150$ s $V$ >>21.96 This work
59.291 1.5m OSN $12\times 150$ s $R$ 22.19 0.18 This work
59.336 1.5m OSN $12\times 150$ s II 23.01 0.28 This work
GRB 240225B (T0= 2460366.34428); GOTO24tz
1.501 GOTO-N $4\times 45$ s $L$ 17.118 0.043 This work
18.90 0.5m HMT $30\times 90$ s $R$ 18.88 0.12 This work
25.29 0.5m HMT $30\times 90$ s $R$ 19.81 0.3 This work
25.914 GOTO-N $4\times 45$ s $L$ 19.694 0.178 This work
26.167 ALFOSC/NOT $3\times 300$ s $r^{\prime}$ 19.159 0.016 This work
28.096 IO:O/LT $2\times 75$ s $g$ 19.971 0.123 This work
28.150 IO:O/LT $2\times 75$ s $r$ 19.513 0.151 This work
28.204 IO:O/LT $2\times 75$ s $i$ 19.442 0.105 This work
28.258 IO:O/LT $2\times 75$ s $z$ 19.281 0.143 This work
39.312 50cm MITSuME-Akeno $79\times 60$ s $g$ 19.78 0.13 Sasada et al. (2024)
39.312 50cm MITSuME-Akeno $79\times 60$ s $R$ 19.53 0.10 Sasada et al. (2024)
39.312 50cm MITSuME-Akeno $79\times 60$ s II >>19.7 Sasada et al. (2024)
41.261 CMOS/AZT-33IK (Mondy) $60\times 60$ s $R$ 19.93 0.07 This work
49.359 IO:O/LT $15\times 120$ s $r$ 20.04 0.05 This work
52.804 3.6m DOT $60$ s $i$ 19.96 0.04 Ror et al. (2024)
57.656 WINTER/Palomar $8\times 120$ s $J$ >>18.8 Mo et al. (2024)
57.656 WINTER/Palomar $8\times 120$ s $Y$ >>18.3 Mo et al. (2024)
60.238 IO:O/LT $15\times 120$ s $r$ 20.444 0.124 This work
68.194 CMOS/AZT-33IK (Mondy) $60\times 60$ s $R$ 20.64 0.11 This work
70.337 Zeiss-1000 (SAO-RAS) $8\times 300$ s $R$ 20.72 0.02 Moskvitin et al. (2024a)
73.160 IO:O/LT $15\times 180$ s $r$ 21.303 0.126 This work
75.762 ALFOSC/NOT $3\times 300$ s $r^{\prime}$ 20.660 0.052 This work
77.280 X-shooter/VLT $60$ s $r^{\prime}$ 20.879 0.031 This work
77.324 X-shooter/VLT $3\times 60$ s $r^{\prime}$ 20.956 0.035 This work
77.410 X-shooter/VLT $3\times 60$ s $g^{\prime}$ 21.055 0.082 This work
77.477 X-shooter/VLT $3\times 60$ s $z^{\prime}$ 20.79 0.129 This work
93.367 Zeiss-1000 (SAO RAS) $8\times 300$ s $R$ 21.32 0.04 Moskvitin et al. (2024b)
113.026 CMOS/AZT-33IK (Mondy) $60\times 60$ s $R$ 21.73 0.19 This work
137.128 CMOS/AZT-33IK (Mondy) $60\times 60$ s $R$ 21.92 0.24 This work
235.750 CMOS/AZT-33IK (Mondy) $90\times 60$ s $R$ >>23.02 This work
259.086 CMOS/AZT-33IK (Mondy) $114\times 60$ s $R$ >>23.92 This work
283.934 CMOS/AZT-33IK (Mondy) $116\times 60$ s $R$ >>23.82 This work
308.135 CMOS/AZT-33IK (Mondy) $120\times 60$ s $R$ >>23.82 This work
379.454 CMOS/AZT-33IK (Mondy) $150\times 60$ s $R$ >>23.72 This work
450.107 CMOS/AZT-33IK (Mondy) $142\times 60$ s $R$ >>23.31 This work
476.372 CMOS/AZT-33IK (Mondy) $150\times 60$ s $R$ >>23.39 This work
GRB 240619A (T0= 2460480.65522); GOTO24cvn; need to add GTC and VLT data
2.532 ATLAS $30$ s $o$ 16.242 0.014 ATLAS FP
4.689 GOTO-S $3\times 90$ s $L$ 17.171 0.170 This work
17.955 GOTO-N $4\times 90$ s $L$ 18.381 0.086 This work
26.705 ATLAS $30$ s $o$ 18.724 0.125 ATLAS FP
48.010 WINTER/1m Palomar $30\times 120$ s $J$ 19.3 Mo et al. (2023)
48.810 UVOT/Swift $81.3$ s $v$ >>18.78 This work
48.684 UVOT/Swift $628$ s whitewhite 20.64 0.10 This work
49.174 UVOT/Swift $1070.9$ s $u$ 20.29 0.14 This work
66.054 ALFOSC/NOT $5\times 300$ s $r^{\prime}$ 19.489 0.030 This work
65.482 ALFOSC/NOT $9\times 200$ s $z^{\prime}$ 20.01 0.04 This work
98.758 PS1-GPC1 $300$ s $i$ 20.86 0.28 TNS, 215607
113.909 ALFOSC/NOT $5\times 300$ s $r^{\prime}$ 20.452 0.037 This work
161.494 ALFOSC/NOT $5\times 300$ s $r^{\prime}$ 20.681 0.051 This work
331.618 X-shooter/VLT $10\times 10$ s $r^{\prime}$ 21.705 0.101 This work
GRB 240910A (T0= 2460563.66718); GOTO24fvl
9.430 GOTO-S $4\times 90$ s $L$ 19.329 0.130 This work
10.559 GOTO-S $4\times 90$ s $L$ 19.879 0.152 This work
10.885 GOTO-S $3\times 90$ s $L$ 20.161 0.243 This work
11.689 GOTO-S $4\times 90$ s $L$ 19.970 0.320 This work
12.014 GOTO-S $4\times 90$ s $L$ 19.740 0.123 This work
36.613 UVOT/Swift $1561.5$ s $u$ 21.59 0.18 This work
36.758 UVOT/Swift $378.7$ s $v$ >>20.08 This work
37.344 ALT/100C $10\times 300$ s $r$ 20.8 0.2 This work
43.504 ALFOSC/NOT $3\times 300$ s $r^{\prime}$ 20.985 0.047 This work
45.428 IO:O/LT $6\times 180$ s $r$ 21.12 0.13 This work
45.908 IO:O/LT $6\times 180$ s $z$ 20.76 0.09 This work
46.148 IO:O/LT $6\times 180$ s $i$ 21.09 0.10 This work
47.066 OSIRIS/GTC $30$ s $r^{\prime}$ 21.156 0.028 This work
57.841 VT/SVOM $R$ 21.50 0.05 SVOM/VT Team et al. (2024)
57.841 VT/SVOM $B$ 22.24 0.07 SVOM/VT Team et al. (2024)
84.022 UVOT/Swift $670.8$ s $u$ >>22.04 This work
84.958 UVOT/Swift $1758.4$ s $v$ >>20.96 This work
91.237 UVOT/Swift $315.7$ s $u$ >>21.53 This work
191.928 UVOT/Swift $3313.0$ s $u$ >>22.95 This work
196.643 UVOT/Swift $768.2$ s $u$ >>22.19 This work
386.321 LBC/LBT $900$ s $z^{\prime}$ >>24.5 This work
386.321 LBC/LBT $900$ s $r^{\prime}$ 25.99 0.35 This work
GRB 240916A (T0= 2460569.557581); GOTO24fzn
7.731 GOTO-S $4\times 90$ s $L$ 17.801 0.055 This work
12.872 ALT/100C $4\times 300$ s $r$ 18.60 0.05 This work
17.064 UVOT/Swift $237$ s whitewhite 20.93 0.18 This work
17.126 UVOT/Swift $188.2$ s $v$ 19.08 0.31 This work
18.339 UVOT/Swift $1185.6$ s $u$ 21.14 0.24 This work
22.750 X-shooter/VLT $19\times 30$ s $r^{\prime}$ 19.437 0.023 This work
42.674 ALFOSC/NOT $3\times 300$ s $r^{\prime}$ 20.433 0.035 This work
46.373 X-shooter/VLT $7\times 30$ s $r^{\prime}$ 20.627 0.033 This work
60.595 ALT/100C $6\times 300$ s $r$ >>19.2 This work
84.684 ALT/100B $8\times 180$ s $r$ >>19.9 This work
101.853 UVOT/Swift $916.3$ s whitewhite >>22.60 This work
156.200 UVOT/Swift $1693.2$ s whitewhite >>22.92 This work
GRB 241002B (T0= 2460585.75993); GOTO24gpc
3.051 GOTO-S $4\times 90$ s $L$ 19.53 0.09 This work
6.206 QHY600 CMOS/40cm LCOGT $500$ s $g$ 20.18 0.3 Torreiro Martínez et al. (2024)
6.348 QHY600 CMOS/40cm LCOGT $500$ s $r$ 19.90 0.25 Torreiro Martínez et al. (2024)
16 1.8m PRIME $J$ 20.0 0.2 Durbak et al. (2024a)
16 1.8m PRIME $H$ 19.6 0.1 Durbak et al. (2024a)
38.986 UVOT/Swift $2619.0$ s $u$ 21.99 0.18 This work
40 1.8m PRIME $J$ 20.6 0.2 Durbak et al. (2024b)
40 1.8m PRIME $H$ 20.3 0.1 Durbak et al. (2024b)
64 1.8m PRIME $H$ 20.8 0.2 Guiffreda et al. (2024)
172.080 UVOT/Swift $4302.1$ s $u$ >>23.17 This work
GRB 241228B (T0= 2460672.67575); GOTO24jmz
0.322 GOTO-N $4\times 90$ s $L$ 14.543 0.007 This work
0.349 GOTO-N $4\times 90$ s $L$ 14.647 0.008 This work
0.377 GOTO-N $4\times 90$ s $L$ 14.713 0.008 This work
0.405 GOTO-N $4\times 90$ s $L$ 14.778 0.009 This work
1.483 GOTO-N $4\times 90$ s $L$ 17.001 0.045 This work
5.420 1m LCOGT $600$ s $r^{\prime}$ 18.62 0.06 Ortega-Casas et al. (2024)
5.899 1m LCOGT $600$ s $i^{\prime}$ 18.31 0.09 Ortega-Casas et al. (2024)
6.095 1m LCOGT $600$ s $g^{\prime}$ 19.40 0.06 Ortega-Casas et al. (2024)
6.609 1m LCOGT $600$ s $z^{\prime}$ 18.25 0.28 Ortega-Casas et al. (2024)
7.602 ATLAS $30$ s $o$ 18.861 0.057 ATLAS FP
8.664 0.7m TRT/SBO $4\times 300$ s $R$ 19.41 0.04 This work
8.794 GOTO-S $4\times 90$ s $L$ 19.702 0.099 This work
9.957 1m LCOGT $600$ s $r^{\prime}$ 19.62 0.05 Ortega-Casas et al. (2024)
10.976 1m LCOGT $1200$ s $B$ 21.35 0.03 Ghosh et al. (2024)
11.362 UVOT/Swift $1778.5$ s $u$ 21.98 0.33 This work
11.628 1.6m Mephisto $300\times 4$ s $u_{m}$ 21.69 0.17 This work
11.628 1.6m Mephisto $300\times 4$ s $g_{m}$ 20.36 0.08 This work
11.629 1.6m Mephisto $300\times 4$ s $i_{m}$ 19.41 0.05 This work
14.175 1.6m Mephisto $300\times 3$ s $v_{m}$ 21.71 0.22 This work
14.175 1.6m Mephisto $300\times 3$ s $r_{m}$ 20.27 0.05 This work
14.212 1.6m Mephisto $300\times 2$ s $i_{m}$ 20.30 0.24 This work
15.739 1.6m Mephisto $300\times 3$ s $v_{m}$ 21.70 0.23 This work
15.739 1.6m Mephisto $300\times 3$ s $r_{m}$ 20.42 0.07 This work
15.740 1.6m Mephisto $300\times 3$ s $i_{m}$ 20.10 0.06 This work
15.991 CMOS/AZT-33IK (Mondy) $30\times 120$ s $R$ 20.49 0.04 This work
18 D50 $24\times 120$ s $r^{\prime}$ 20.4 0.1 Strobl & Jelinek (2024)
18.469 1.6m Mephisto $300\times 3$ s $v_{m}$ 21.38 0.23 This work
18.469 1.6m Mephisto $300\times 3$ s $r_{m}$ 20.54 0.08 This work
18.469 1.6m Mephisto $300\times 3$ s $i_{m}$ 20.35 0.10 This work
20.327 ALFOSC/NOT $2\times 150$ s $r^{\prime}$ 20.560 0.061 This work
25.591 X-shooter/VLT $10$ s $r^{\prime}$ 20.626 0.029 This work
25.640 X-shooter/VLT $3\times 60$ s $r^{\prime}$ 20.617 0.032 This work
25.688 X-shooter/VLT $3\times 60$ s $g^{\prime}$ 21.448 0.011 This work
25.775 X-shooter/VLT $3\times 60$ s $z^{\prime}$ 20.267 0.023 This work
33.933 GOTO-S $4\times 45$ s $L$ >>20.52 This work
39.066 1.6m Mephisto $300\times 4$ s $v_{m}$ >>22.85 This work
39.066 1.6m Mephisto $300\times 4$ s $r_{m}$ 21.02 0.09 This work
39.068 1.6m Mephisto $300\times 4$ s $i_{m}$ 20.72 0.10 This work
41.972 1.6m Mephisto $300\times 5$ s $v_{m}$ >>22.80 This work
41.972 1.6m Mephisto $300\times 5$ s $r_{m}$ 21.24 0.11 This work
41.972 1.6m Mephisto $300\times 3$ s $i_{m}$ 20.94 0.14 This work
42.175 Zeiss-1000 (SAO RAS) $12\times 300$ s $R$ 21.18 0.06 Moskvitin et al. (2024c)
61.619 1.3m DFOT $24\times 300$ s $R$ 21.44 0.05 Ror et al. (2025)
63.023 1.6m Mephisto $300\times 4$ s $v_{m}$ >>22.42 This work
63.023 1.6m Mephisto $300\times 4$ s $r_{m}$ 21.87 0.15 This work
63.025 1.6m Mephisto $300\times 4$ s $i_{m}$ 21.50 0.14 This work
66.200 1.6m Mephisto $300\times 4$ s $v_{m}$ >>22.74 This work
66.200 1.6m Mephisto $300\times 4$ s $r_{m}$ 22.49 0.19 This work
66.203 1.6m Mephisto $300\times 4$ s $i_{m}$ 21.92 0.19 This work
87.254 1.6m Mephisto $300\times 5$ s $v_{m}$ >>22.81 This work
87.254 1.6m Mephisto $300\times 5$ s $r_{m}$ 22.70 0.20 This work
87.254 1.6m Mephisto $300\times 5$ s $i_{m}$ >>23.40 This work
204.161 UVOT/Swift $4273.9$ s $u$ >>22.59 This work
Table A2: Log of radio follow-up observations. The flux density errors are 1 sigma and the upper limits correspond to three times the image RMS.
   Observation Date
   Telescope
   Time Post-burst
   Frequency
   Flux Density
   Source
   GRB 240122A
   2024-01-24 12:01:00.0 UT    ATCA    2.06    5.5    <159.2    Anderson et al. (2024b)
   ATCA    9.0    160.0 $\pm$ 20.0    Anderson et al. (2024b)
   2024-01-26 06:39:24.9 UT    ATCA    3.84    5.5    <1555.0    This work
   ATCA    9.0    <100.5    This work
   ATCA    16.7    <113.1    This work
   ATCA    21.2    <333.0    This work
   2024-01-28 10:37:34.9 UT    ATCA    6.01    5.5    <174.0    This work
   ATCA    9.0    66.3 $\pm$ 19.7    This work
   ATCA    16.7    <55.5    This work
   ATCA    21.2    <111.3    This work
   2024-02-12 03:03:54.9 UT    ATCA    20.69    5.5    <45.0    This work
   ATCA    9.0    <33.6    This work
   GRB 240619A
   2024-06-22 13:45:34 UT    AMI-LA    3.42    15.5    $1500\pm 60$    Rhodes et al. (2024b)
   GRB 240910A
   2024-09-13 06:34:25.6 UT    VLA    3.11    6    $137\pm 10$    Giarratana et al. (2024a)
   2024-09-13 06:56:22.1 UT    VLA    3.12    10    $114\pm 9$    Giarratana et al. (2024a)
   2024-09-13 06:16:03.7 UT    VLA    3.09    15    $86\pm 10$    Giarratana et al. (2024a)
   2024-09-19 06:33:15.8 UT    VLA    9.10    6    $58\pm 7$    This work
   2024-09-19 06:54:57.2 UT    VLA    9.12    10    $47\pm 8$    This work
   2024-09-19 06:14:53.8 UT    VLA    9.09    15    $71\pm 10$    This work
   2024-10-01 11:12:12.0 UT    VLA    21.30    6    $22\pm 6$    This work
   2024-10-01 11:34:09.0 UT    VLA    21.31    10    <24    This work
   2024-10-01 10:53:51.0 UT    VLA    21.29    15    <27    This work
   2024-10-26 09:15:46.0 UT    VLA    46.22    6    <18    This work
   2024-10-26 09:38:12.0 UT    VLA    46.23    10    <27    This work
   2024-10-26 08:55:00.0 UT    VLA    46.20    15    <18    This work
   GRB 240916A
   2024-09-18 00:43:57 UT    VLA    1.97    6    $35\pm 8$    Giarratana et al. (2024b)
   VLA    10    $44\pm 8$    Giarratana et al. (2024b)
   VLA    15    $135\pm 8$    Giarratana et al. (2024b)
Figure A1: X-shooter spectra of GRB 240122A at $z=3.1634\pm 0.0003$, observed at 16.130 hr post-trigger. The three panels show the spectra from the UVB, VIS, and NIR arms (top to bottom). Each panel displays the 2D spectrum (upper sub-panel) and the corresponding 1D extracted spectrum in black with the error spectrum in grey (lower sub-panel). Absorption lines identified at the redshift of the GRB are marked in blue and labelled, while grey-shaded regions indicate telluric absorption. The 1D spectra have been smoothed with a Savitzky–Golay filter to enhance the visibility of spectral features. The same colour scheme is adopted for all spectra presented throughout this work.
Figure A2: GTC/OSIRIS spectrum of GRB 240122A at $z=3.1634\pm 0.0003$, obtained 12.857 hr after the trigger. The upper panel displays the 2D spectrum, and the lower panel shows the extracted 1D spectrum in black with its error spectrum in grey. Absorption features at the GRB redshift are highlighted in blue and labelled, while grey-shaded regions mark telluric absorption. The 1D spectra were smoothed using a Savitzky–Golay filter to improve the visibility of spectral features.
Figure A3: VLT/X-shooter spectra of GRB 240225B at $z=0.9462\pm 0.0002$, taken at 3.237 d post-trigger. Absorption lines identified at the GRB redshift are marked in blue and labelled accordingly, while emission lines are indicated by black dotted vertical lines. A Savitzky-Golay filter was applied to the 1D spectra to enhance the clarity of spectral features.
Figure A4: VLT/X-shooter spectra of GRB 240619A at $z=0.3960\pm 0.0001$, observed at 13.826 d post-trigger. Emission lines are indicated by black dotted vertical lines and labelled accordingly. To highlight spectral features, the 1D spectra were smoothed with a Savitzky–Golay filter.
Figure A5: GTC/OSIRIS spectrum of GRB 240910A at $z=1.4605\pm 0.0007$ at 1.964 days post-trigger. The upper panel presents the 2D spectrum, while the lower panel shows the extracted 1D spectrum in black with its associated error in grey. Absorption features at the GRB redshift are indicated in blue and labelled, and grey-shaded regions denote telluric absorption. For improved feature visibility, the 1D spectra were smoothed using a Savitzky-Golay filter.
Figure A6: VLT/X-shooter spectra of GRB 240916A at $z=2.6100\pm 0.0002$, taken at 22.962 hr post-trigger. Absorption lines identified at the redshift of the GRB are marked in blue and labelled accordingly. The 1D spectra were filtered with a Savitzky–Golay function to enhance the prominence of spectral features.
Figure A7: VLT/X-shooter spectra of GRB 241228B at $z=2.6745\pm 0.0004$, obtained at 25.975 hr post-trigger. Absorption lines identified at the redshift of the GRB are marked in blue and labelled accordingly. The 1D spectra have been smoothed with a Savitzky–Golay filter to enhance the visibility of spectral features.
Figure A8: MAXI/GSC lightcurves for GRBs in our sample (240122A, 240225B) are shown. Data are binned to 1-sec resolution.
Figure A9: MAXI fitted spectra for GRBs 240122A and 240225B are shown. Models shown are the best-fit models as described in Table 4. GRB 241228B is poorly constrained in high energies, so the spectrum shown is zoomed in to show the behaviour at lower energies.
Figure A10: Fermi/GBM lightcurves for Fermi GRBs in our sample are shown. TTE data are binned to 0.5s resolution.
Figure A11: Fermi/GBM fitted spectra for Fermi GRBs in our sample are shown. Models shown are the best-fit models as described in Table 4. GRB 241228B is poorly constrained in high energies, so the spectrum shown is zoomed in to show the behaviour at lower energies.
Figure A12: Swift/XRT count-rate spectra and best-fitting X-ray continuum models for the GRBs in our sample (GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B). In each panel, the top plot displays observed spectral data with errors (black points) in the 0.3–10 keV range fitted with an absorbed power-law model (red line), using Cash statistics. The bottom plot shows the ratio of observed data to folded model predictions, used to assess the goodness of fit. These spectra were used to derive photon indices and absorption column densities ($N_{\mathrm{H}}$), contributing to the multi-wavelength characterisation of the afterglows.
Table A3: Absorption features identified in the afterglow spectrum of GRB 240122A ($z=3.1634\pm 0.0003$), based on observations with VLT/X-shooter and GTC/OSIRIS. The UVB, VIS, and NIR designations refer to the respective arms of the VLT/X-shooter spectrograph.
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
3796.2 911.8 Ly limit abs UVB
[…]
5247.9 1260.4 Si ii abs UVB
5556.5 1334.5 C ii abs UVB
5803.1 1393.8 Si iv abs VIS
5840.6 1402.8 Si iv abs VIS
6356.6 1526.7 Si ii abs VIS
6446.1 1548.2 C iv abs VIS
6456.8 1550.8 C iv abs VIS
6956.5 1670.8 Al ii abs VIS
11642.9 2796.4 Mg ii abs NIR
11672.8 2803.5 Mg ii abs NIR
Table A4: Absorption and emission features identified in the afterglow spectrum of GRB 240225B ($z=0.9462\pm 0.0002$), obtained with VLT/X-shooter.
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
3610.2 1854.7 Al iii abs UVB
3625.9 1862.8 Al iii abs UVB
4002.5 2056.3 Cr ii abs UVB
4379.4 2249.9 Fe ii abs UVB
4563.0 2344.2 Fe ii abs UVB
4621.9 2374.5 Fe ii abs UVB
4638.1 2382.8 Fe ii abs UVB
4650.9 2389.4 Fe ii abs UVB
4664.5 2396.4 Fe ii abs UVB
4671.6 2400.0 Fe ii abs UVB
5015.9 2576.9 Mn ii abs UVB
5034.9 2586.6 Fe ii abs UVB
5050.2 2594.5 Mn ii abs UVB
5061.2 2600.2 Fe ii abs UVB
5443.1 2796.4 Mg ii abs UVB
5457.1 2803.5 Mg ii abs UVB
5553.4 2853.0 Mg i abs UVB
7254.8 3727.1 [O ii] em VIS
7260.2 3729.9 [O ii] em VIS
[…]
7726.8 3969.6 Ca ii abs VIS
9748.5 5008.2 [O iii] em VIS
Table A5: Absorption and emission features identified in the afterglow spectrum of GRB 240619A ($z=0.3960\pm 0.0001$), obtained with VLT/X-shooter.
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
5203.0 3727.1 [O ii] em UVB
5206.9 3729.9 [O ii] em UVB
5402.7 3870.2 [Ne iii] em UVB
[…]
6924.6a 4960.3 [O iii] em VIS
6991.5 5008.2 [O iii] em VIS
[…]
Table A6: Absorption features identified in the afterglow spectrum of GRB 240910A ($z=1.4605\pm 0.0007$), obtained with GTC/OSIRIS. Here, a marks lines affected by telluric absorption.
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type
3755.5 1526.7 Si ii abs
3808.4 1548.2 C iv abs
3956.8 1608.5 Fe ii abs
3967.3 1612.8 Fe ii abs
4110.0 1670.8 Al ii abs
4447.5 1808.0 Si ii abs
4562.4 1854.7 Al iii abs
4984.5 2026.3 Cr ii abs
5058.3 2056.3 Cr ii abs
5082.6 2066.2 Cr ii abs
5328.6 2166.2 Ni ii abs
5376.4 2185.6 Mn i abs
5561.3 2260.8 Fe ii abs
5699.1 2316.8 Ni ii abs
5766.5 2344.2 Fe ii abs
5841.0 2374.5 Fe ii abs
5861.4 2382.8 Fe ii abs
5894.7 2396.3 Fe ii abs
6338.9a 2576.9 Mn ii abs
6363.0a 2586.7 Fe ii abs
6396.2a 2600.2 Fe ii abs
6878.6 2796.3 Mg ii abs
6896.3 2803.5 Mg ii abs
7018.0 2852.7 Mg i abs
Table A7: Absorption features identified in the afterglow spectrum of GRB 240916A ($z=2.6100\pm 0.0002$), obtained with VLT/X-shooter. Superscript symbols indicate excited-state transitions (, ∗∗), while a marks lines affected by telluric absorption.
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
[…]
4513.2 1250.2 S ii abs UVB
4524.8 1253.4 S ii abs UVB
4546.4 1259.4 S ii abs UVB
4566.1 1264.8 Si ii abs UVB
4700.8 1302.2 O i abs UVB
4708.8 1304.4 Si ii abs UVB
4710.5 1304.9 O i abs UVB
4714.8 1306.0 O i∗∗ abs UVB
4726.5 1309.3 Si ii abs UVB
4817.7 1334.5 C ii abs UVB
4821.9 1335.7 C ii abs UVB
5031.5 1393.8 Si iv abs UVB
5064.0 1402.8 Si iv abs UVB
5511.4 1526.7 Si ii abs UVB
5535.7 1533.4 Si ii abs UVB
5589.0 1548.2 C iv abs VIS
5598.3 1550.8 C iv abs VIS
5806.5 1608.5 Fe ii abs VIS
5822.2 1612.8 Fe ii abs VIS
6031.5 1670.8 Al ii abs VIS
6144.2 1702.0 Fe ii abs VIS
6287.0 1741.6 Ni ii abs VIS
6526.9 1808.0 Si ii abs VIS
6695.5 1854.7 Al iii abs VIS
6724.7 1862.8 Al iii abs VIS
7315.6 2026.5 Mg i abs VIS
7314.4 2026.1 Zn ii abs VIS
7314.9 2026.3 Cr ii abs VIS
7423.1 2056.3 Cr ii abs VIS
7444.5 2062.2 Cr ii abs VIS
7445.4 2062.4 Zn ii abs VIS
7459.0 2066.2 Cr ii abs VIS
7820.0 2166.2 Ni ii abs VIS
8004.1 2217.2 Ni ii abs VIS
8027.2 2223.6 Ni ii abs VIS
8122.1 2249.9 Fe ii abs VIS
8161.5 2260.8 Fe ii abs VIS
8363.3 2316.7 Ni ii abs VIS
8404.4 2328.1 Fe ii abs VIS
8423.9 2333.5 Fe ii abs VIS
8442.7 2338.7 Fe ii abs VIS
8462.6 2344.2 Fe ii abs VIS
8479.9 2349.0 Fe ii abs VIS
8539.8 2365.6 Fe ii abs VIS
8571.8 2374.5 Fe ii abs VIS
8601.8 2382.8 Fe ii abs VIS
8625.6 2389.4 Fe ii abs VIS
8650.8 2396.4 Fe ii abs VIS
8664.0 2400.0 Fe ii abs VIS
8684.2 2405.6 Fe ii abs VIS
8690.7 2407.4 Fe ii abs VIS
8704.8 2411.3 Fe ii abs VIS
8714.9 2414.1 Fe ii abs VIS
9302.5a 2576.9 Mn ii abs VIS
9337.8a 2586.6 Fe ii abs VIS
9386.6a 2600.2 Fe ii abs VIS
9409.3a 2606.5 Mn ii abs VIS
9414.5a 2607.9 Fe ii abs VIS
9431.8a 2612.7 Fe ii abs VIS
9452.4a 2618.4 Fe ii abs VIS
9491.1a 2629.1 Fe ii abs VIS
9501.3a 2631.9 Fe ii abs VIS
10094.8 2796.4 Mg ii abs VIS
10120.7 2803.5 Mg ii abs VIS
Table A8: Absorption and emission features identified in the afterglow spectrum of GRB 241228B ($z=2.6745\pm 0.0004$), obtained with VLT/X-shooter. Superscript symbols denote excited-state transitions (, ∗∗, etc.), while a marks features affected by telluric absorption.
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
[…]
4551.8 1238.8 N v abs UVB
4566.5 1242.8 N v abs UVB
4593.7 1250.2 S ii abs UVB
4605.5 1253.4 S ii abs UVB
4627.5 1259.4 S ii abs UVB
4631.2 1260.4 Si ii abs UVB
4647.6 1264.8 Si ii abs UVB
4784.8 1302.2 O i abs UVB
4792.9 1304.4 Si ii abs UVB
4794.7 1304.9 O i abs UVB
4798.8 1306.0 O i∗∗ abs UVB
4810.9 1309.3 Si ii abs UVB
4839.9 1317.2 Ni ii abs UVB
4903.1 1334.4 C ii abs UVB
4907.9 1335.7 C ii abs UVB
5034.3 1370.1 Ni ii abs UVB
5121.2 1393.8 Si iv abs UVB
5154.3 1402.8 Si iv abs UVB
5345.5 1454.8 Ni ii abs UVB
5391.4 1467.3 Ni ii abs UVB
5393.3 1467.8 Ni ii abs UVB
5609.7 1526.7 Si ii abs VIS
5634.3 1533.4 Si ii abs VIS
5688.7 1548.2 C iv abs VIS
5698.3 1550.8 C iv abs VIS
5728.4 1559.0 Fe ii∗∗∗∗∗ abs VIS
5757.0 1566.8 Fe ii abs VIS
5910.1 1608.5 Fe ii abs VIS
5920.2 1611.2 Fe ii abs VIS
5926.1 1612.8 Fe ii abs VIS
6016.5 1637.4 Fe ii abs VIS
6139.0 1670.8 Al ii abs VIS
6253.8 1702.0 Fe ii abs VIS
6281.8 1709.6 Ni ii abs VIS
6399.3 1741.6 Ni ii abs VIS
6437.2 1751.9 Ni ii abs VIS
6643.3 1808.0 Si ii abs VIS
6676.0 1816.9 Si ii abs VIS
6678.2 1817.5 Si ii abs VIS
6716.4 1827.9 Mg i abs VIS
6814.9 1854.7 Al iii abs VIS
6844.5 1862.8 Al iii abs VIS
7444.7 2026.1 Zn ii abs VIS
7445.4 2026.3 Cr ii abs VIS
7446.2 2026.5 Mg i abs VIS
7555.7 2056.3 Cr ii abs VIS
7577.3 2062.2 Cr ii abs VIS
7579.0a 2062.7 Zn ii abs VIS
7592.0a 2066.2 Cr ii abs VIS
7959.5 2166.2 Ni ii abs VIS
8146.9 2217.2 Ni ii abs VIS
8170.4 2223.6 Ni ii abs VIS
8267.0 2249.9 Fe ii abs VIS
8307.1 2260.8 Fe ii abs VIS
8512.8 2316.8 Ni ii abs VIS
8554.4 2328.1 Fe ii abs VIS
8574.2 2333.5 Fe ii abs VIS
8593.3 2338.7 Fe ii abs VIS
8613.5 2344.2 Fe ii abs VIS
Table A9: GRB 241228B (continued)
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
8616.5 2345.0 Fe ii abs VIS
8631.2 2349.0 Fe ii abs VIS
8670.8 2359.8 Fe ii abs VIS
8692.2 2365.6 Fe ii abs VIS
8724.9 2374.5 Fe ii abs VIS
8750.6 2381.5 Fe ii abs VIS
8755.4 2382.8 Fe ii abs VIS
8779.6 2389.4 Fe ii abs VIS
8805.0 2396.3 Fe ii abs VIS
8818.6 2400.0 Fe ii abs VIS
8845.8 2407.4 Fe ii abs VIS
8859.9 2411.2 Fe ii abs VIS
8861.9 2411.8 Fe ii abs VIS
9468.6a 2576.9 Mn ii abs VIS
9504.4a 2586.7 Fe ii abs VIS
9533.2a 2594.5 Mn ii abs VIS
9554.2a 2600.2 Fe ii abs VIS
9577.3a 2606.5 Mn ii abs VIS
9582.5a 2607.9 Fe ii abs VIS
9600.1a 2612.7 Fe ii abs VIS
9607.1a 2614.6 Fe ii abs VIS
9621.0a 2618.4 Fe ii abs VIS
9636.1a 2622.5 Fe ii abs VIS
9650.8a 2626.5 Fe ii abs VIS
9660.4a 2629.1 Fe ii abs VIS
10274.9 2796.3 Mg ii abs NIR
10301.2 2803.5 Mg ii abs NIR
10483.1 2853.0 Mg i abs NIR
14585.9 3969.6 Ca ii abs NIR
15535.0 4227.9 Ca i abs NIR
Affiliations
1Department of Physics, Royal Holloway - University of London, Egham, TW20 0EX, U.K.
2Department of Physics, University of Warwick, Gibbet Hill Road, Coventry CV4 7AL, UK.
3School of Physics and Astronomy, University of Birmingham, Edgbaston, Birmingham, B15 2TT, UK.
4Institute for Gravitational Wave Astronomy, University of Birmingham, Birmingham, B15 2TT, UK.
5Aix Marseille Univ., CNRS, CNES, LAM, Marseille, France.
6School of Physics & Astronomy, Monash University, Clayton VIC 3800, Australia.
7Université Paris-Saclay, Université Paris Cité, CEA, CNRS, AIM, 91191, Gif-sur-Yvette, France.
8School of Physics and Astronomy, University of Leicester, University Road, Leicester, LE1 7RH, United Kingdom.
9Sydney Institute for Astronomy, School of Physics, The University of Sydney, NSW 2006, Australia.
10Centre of Excellence for Gravitational Wave Discovery (OzGrav), Hawthorn, VIC 3122, Australia.
11CSIRO Space and Astronomy, PO Box 76, Epping, NSW 1710, Australia.
12Department of Astrophysics/IMAPP, Radboud University, 6525 AJ Nijmegen, The Netherlands.
13Niels Bohr Institute, University of Copenhagen, Jagtvej 155, 2200, Copenhagen N, Denmark
14The Cosmic Dawn Centre (DAWN), Denmark
15Astrophysics Research Cluster, School of Mathematical and Physical Sciences, University of Sheffield, Sheffield, S3 7RH, UK.
16Research Software Engineering, University of Sheffield, Sheffield, S1 4DP, UK.
17INAF Osservatorio Astronomico di Brera, Via E. Bianchi 46, I-23807 Merate, Italy.
18Department of Physical Science, Aoyama Gakuin University, 5-10-1 Fuchinobe, Chuo-ku, Sagamihara, Kanagawa 252-5258, Japan.
19South-Western Institute for Astronomy Research (SWIFAR), Yunnan University, Kunming, Yunnan 650500, People’s Republic of China.
20Yunnan Key Laboratory of Survey Science, Yunnan University, Kunming, Yunnan 650500, People’s Republic of China.
21National Astronomical Observatories, Chinese Academy of Sciences, Beijing 100012, China.
22Guangxi Key Laboratory for Relativistic Astrophysics, School of Physical Science and Technology, Guangxi University, Nanning 530004, China.
23Armagh Observatory and Planetarium, College Hill, Armagh, BT61 9DG, N Ireland, UK.
24Institute for Globally Distributed Open Research and Education (IGDORE).
25Instituto de Astrofísica de Canarias, E-38205 La Laguna, Tenerife, Spain.
26National Astronomical Research Institute of Thailand, 260 Moo 4, T. Donkaew, A. Maerim, Chiangmai, 50180 Thailand.
27Department of Physics & Astronomy, University of Turku, Vesilinnantie 5, Turku, FI-20014, Finland.
28Jodrell Bank Centre for Astrophysics, Department of Physics and Astronomy, The University of Manchester, Manchester, M13 9PL, UK.
29Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK.
30School of Physics, University College Cork, Cork, Ireland.
31Centre for Astrophysics Research, University of Hertfordshire, College Lane, Hatfield, AL10 9AB, UK.
32School of Physics and Centre for Space Research, University College Dublin, Belfield, Dublin 4, Ireland.
33National Research University ‘Higher School of Economics’, Faculty of Physics, Myasnitskaya ul. 20, Moscow 101000, Russia.
34Space Research Institute of the Russian Academy of Sciences, Profsoyuznaya ul. 84/32, Moscow 117997, Russia.
35Centro Astronómico Hispano en Andalucía, Observatorio de Calar Alto, Sierra de los Filabres, Gérgal, Almería, 04550, Spain
36Departament d’Astronomía i Astrofísica, Universitat de València, 46100 Burjassot, Spain
37Observatori Astronòmic, Universitat de València, 46980 Paterna, Spain
38International Centre for Radio Astronomy Research, Curtin University, GPO Box U1987, Perth, WA 6845, Australia.
39 Astrophysics Research Institute, Liverpool John Moores University, 146 Brownlow Hill, Liverpool L3 5RF, UK
40Instituto de Astrofísica de Andalucía (IAA-CSIC), Glorieta de la Astronomía s/n, E-18008, Granada, Spain.
41Ingeniería de Sistemas y Automática, Universidad de Málaga, Unidad Asociada al CSIC porel IAA, Escuela de Ingenierías Industriales, Arquitecto Francisco Peñalosa, 6, Campanillas, 29071 Málaga, Spain.
42Millennium Institute of Astrophysics (MAS), Nuncio Monsenor Sòtero Sanz 100, Providencia, Santiago RM, Chile.
43University of Messina, MIFT Department, Via F.S. D’Alcontres 31, 98166, Messina, Italy
44Space Science Data Center (SSDC) - Agenzia Spaziale Italiana (ASI), Via del Politecnico snc, I-00133 Roma, Italy
45Department of Astronomy, School of Physics, Huazhong University of Science and Technology, Wuhan, 430074, China
46Clemson University, Department of Physics and Astronomy, Clemson, SC 29634-0978, USA
47Osservatorio Astronomico di Capodimonte, INAF, Salita Moiariello 16, Napoli, 80131, Italy
48Institute of Solar-Terrestrial Physics, Russian Academy of Sciences, Siberian Branch, Irkutsk 664033, Russia.
49GRANTECAN S.A., Cuesta de San José s/n, E-38712 Breña Baja, La Palma, Spain.
50INAF - Osservatorio di Astrofisica e Scienza dello Spazio, via Piero Gobetti 93/3, 40129 Bologna, Italy
51GEPI, Observatoire de Paris, Universit´e PSL, CNRS, 5 Place Jules Janssen, Meudon, 92190, France
52Institute of Physics and Technology, Institutskiy Pereulok, 9, Dolgoprudny 141701, Russia.
53Astronomical Institute Anton Pannekoek, University of Amsterdam, PO Box 94249, NL-1090 GE Amsterdam, the Netherlands
54Materials Science and Applied Mathematics, Malmö University, SE-205 06 Malmö, Sweden
55Nordic Optical Telescope, Rambla José Ana Fernández Pérez 7, ES-38711 Breña Baja, Spain
56E. Kharadze Georgian National Astrophysical Observatory, Mt. Kanobili, Abastumani 0301, Adigeni, Georgia.
57Department of Physics, George Washington University, 725 21st St NW, Washington, DC 20052, USA.
58LUX, Observatoire de Paris, Université PSL, CNRS, Sorbonne Université, Meudon, 92190, France